### v1

# Preprint no. NJU-INP 105-25 Exclusive photoproduction of light and heavy vector mesons: thresholds to very high energies

Abstract
A reaction model for $\gamma+p\to V+p$, $V=\rho^{0},\phi,J/\psi,\Upsilon$, which exposes the quark-antiquark content of the photon in making the transition $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$, where $\mathpzc q$} depends on $V$, and couples the intermediate ${\mathpzc q}\bar{\mathpzc q}$ system to the proton’s valence quarks via Pomeron ($\mathbb{P}$) exchange, is used to deliver a unified description of available data – both differential and total cross sections – from near threshold to very high energies, $W$, for all the $V$-mesons. For the $\Upsilon$, this means $10\lesssim W/{\rm GeV}\lesssim 2\,000$. Also provided are predictions for the power-law exponents that are empirically used to characterise the large-$W$ behaviour of the total cross sections and slope parameters characterising the near-threshold differential cross sections. Appealing to notions of vector meson dominance, the latter have been interpreted as vector-meson–proton scattering lengths. The body of results indicate that it is premature to link any $\gamma+p\to V+p$ data with, for instance, in-proton gluon distributions, the quantum chromodynamics trace anomaly, or pentaquark production. Further developments in reaction theory and higher precision data are required before the validity of any such links can be assessed.
journal: Eur. Phys. J. C
1 Introduction
Exclusive vector-meson photoproduction,
$\gamma+p\rightarrow V+p\,,\quad\text{with\,\,}V=\rho,\phi,J/\psi,\Upsilon,$ (1)
provides a clean environment to study diffractive scattering from the proton and thereby obtain information about the reaction mechanism and both hadrons involved in the process.
Understanding the energy dependence of these cross sections has long challenged theory. Contrary to perturbative quantum chromodynamics (QCD) expectations, which suggest that total cross sections should fall as the center‑of‑mass energy-squared, $s=W^{2}$, increases, experiments show that such photoproduction cross sections actually rise with $s$ Baksay et al. (1978). This issue was analysed, e.g., in Ref. Foldy and Peierls (1963), with the conclusion that the mechanism responsible should correspond to exchange of something that has isospin zero and is even under charge conjugation. Interactions of this kind can be expressed in Regge phenomenology via Pomeron, $\mathbb{P}$, exchange, with a trajectory whose intercept $\alpha_{\mathbb{P}}(0)>1$. The QCD character of the Pomeron is widely discussed Donnachie et al. (2004); Gribov (2007). It is typically supposed to represent the collective exchange of a family of colourless, crossing-even states, with two-gluon exchange being the leading contributor Forshaw and Ross (2011). Connections with glueball physics are also discussed Meyer and Teper (2005); Boschi-Filho et al. (2006).
Figure 1: Reaction model for $\gamma+p\to V+p$, referred to hereafter as $\mathbb{P}$-dyn. The quark-antiquark, ${\mathpzc q}\bar{\mathpzc q}$, component of the dressed-photon is probed by Pomeron exchange with the proton, producing an on-shell vector meson, $V$. Referring to Eq. (4): the green rectangle describes Pomeron exchange, $\Gamma^{\mathbb{P}}$, including its couplings to the proton and quark – see Sect. $2.$ Further, the solid black curve is $S_{q}$, the dressed $q$ quark propagator; the shaded lighter-blue circle is $\Gamma_{\mu}^{\gamma}$, the dressed-$\gamma q\bar{q}$ vertex; and the shaded red circle is $\Gamma_{\nu}$, the $V$-meson Bethe-Salpeter amplitude – see Sect. 3. Kinematics: $s=W^{2}=-(p_{2}^{2}+P^{2})$; $t=-(p_{2}-p_{1})^{2}$.
With the high-energy behaviour of $\gamma+p\rightarrow V+p$ described by $\mathbb{P}$ exchange, it is natural to seek an explanation for the cross section at low-energies and determine whether a unified approach to the entire $s$-domain is possible. Both these issues were addressed in Ref. Pichowsky and Lee (1997), which explored the reaction mechanism depicted in Fig. 1 using phenomenological inputs for the required elements. This approach sees the ${\mathpzc q}\bar{\mathpzc q}$ component of the photon, where $\mathpzc q$ is chosen from the set $\{{\mathpzc q}=u/d,s,c,b\}$, depending on $V$, exposed by interaction with the proton via $\mathbb{P}$-exchange and then transformed into the on-shell vector meson when sufficient reaction energy is available, i.e., once the threshold energy is exceeded:
$W_{\rm th}^{V}=m_{p}+m_{V}\,.$ (2)
Here $m_{p}=0.939\,$GeV is the proton mass and $m_{V}$ is that of the vector meson produced.
The $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element in Fig. 1 can be calculated using QCD propagators and vertices, which may be obtained using either continuum or lattice Schwinger function methods (C[L]SMs). There are three nonperturbative, matrix-valued Schwinger functions; and in the Ref. Tang et al. (2024) study of $\gamma+p\to J/\psi+p$, they were calculated using CSMs Eichmann et al. (2016); Burkert and Roberts (2019); Roberts et al. (2021); Binosi (2022); Ding et al. (2023); Ferreira and Papavassiliou (2023); Achenbach et al. (2025). The analysis delivered parameter-free predictions for both differential and total cross sections from near threshold to very high energies that are in agreement with much available data.
Building further upon Refs. Pichowsky and Lee (1997); Tang et al. (2024), it is relevant to determine whether the reaction mechanism indicated in Fig. 1, with elements calculated using CSMs, can deliver realistic, parameter-free predictions for photoproduction reactions involving other vector mesons, both heavier and lighter than the $J/\psi$. The associated comparisons with results from decades of measurements of $\rho$, $\phi$, $J/\psi$, $\Upsilon$ photoproduction reactions, both differential and total cross sections, over a huge range of energies, would provide valuable information on $\mathbb{P}$-exchange universality and other features of photoproduction processes. Notably, spurred by interest in the proton’s gluon content and possible pentaquark states Chen et al. (2020); Anderle et al. (2021); Abdul Khalek et al. (2022), the experimental data set has recently been expanded by experiments at Jefferson Lab (Adhikari et al., 2023, GlueX) and CERN (Aaij et al., 2025, LHCb).
This discussion is arranged as follows. Section 2 presents the reaction model and our Pomeron exchange framework for vector meson photoproduction. It is followed, in Sect. 3, by an explanation of the CSM calculation of the Schwinger functions required to obtain the $\gamma+p\to V+p$ matrix element sketched in Fig. 1. That matrix element enables completion of the cross section calculations. Section 4 explains how one may infer phenomenological vector-meson–proton scattering lengths from near-threshold differential cross section data using notions of vector meson dominance (VMD). Whilst not rigorous, these scattering lengths can be useful in constraining model studies. They even become objective comparative measures after reinterpretation as a near-threshold slope parameter for each cross section. A large collection of numerical results is reported in Sect. 5 along with relevant data comparisons. That discussion continues in Sect. 6, which analyses the large-$W$ behaviour of photoproduction total cross sections. Numerical results for the near-threshold slope parameter (VMD-based VpV-p scattering length) are reported in Sect. 7. Section 8 provides a summary and perspective.
2 Reaction model and Pomeron exchange
2.1 Dynamical photon to meson transition
The image in Fig. 1 corresponds to the following matrix element Pichowsky and Lee (1997); Donnachie and Landshoff (1984, 1987):
${\cal I}_{\mu}^{V}(W,t)$ $=\langle V(P;\lambda)p(p_{2})|\bar{\mathpzc q}\gamma_{\mu}{\mathpzc q}|p(p_{1})\rangle$
$=2t_{\mu\alpha\nu}^{V}(q,P)\epsilon_{\nu}^{\lambda}(P)\bar{u}(p_{2})\tilde{G}_{\alpha}(w^{2},t)u(p_{1})\,$, (3)
where $\epsilon_{\nu}^{\lambda}(P)$ is the $V$-meson polarisation vector, $u(p_{1})$, $\bar{u}(p_{2})$ are spinors for the incoming and outgoing proton, $\tilde{G}_{\alpha}(w^{2},t)$ represents the quark-Pomeron-proton interaction; and the factor “2” expresses the equivalence between striking the upper and lower valence quark lines in the quark loop. The invariant mass of the $\gamma p$ system is $s=W^{2}=-(q+p_{1})^{2}$ and the momentum transfer is $t=-(p_{1}-p_{2})^{2}$.
The exposed quark loop integration in Fig. 1 receives most support on the neighbourhood $k^{2}\simeq 0$, i.e., on the domain of near-zero relative momentum within the final state meson, whereat the meson’s Bethe-Salpeter amplitude is peaked. One may therefore write $w^{2}=-(q-P/2+p_{1})^{2}$. In this case, the essentially dynamical component of the reaction model, viz. the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element, can be written as follows:
$t_{\mu\alpha\nu}^{V}(q,P)=e_{V}N_{c}{\rm tr}_{D}\int\frac{d^{4}k}{(2\pi)^{4}}S_{\mathpzc q}(k_{-})\Gamma_{\mu}^{\gamma}(k_{-},k_{-}+q)$
$\quad\times S_{\mathpzc q}(k_{-}+q)\Gamma^{\mathbb{P}}_{\alpha}(k_{-}+q,k_{+})S_{\mathpzc q}(k_{+})\Gamma_{\nu}^{V}(k_{+},k_{-})\,$, (4)
where $N_{c}=3$; $e_{\rho^{0},\phi,J/\psi,\Upsilon}=e\{1/\surd 2,1/3,2/3,1/3\}$, the fine-structure constant $\alpha=e^{2}/[4\pi]$; the trace is over spinor indices; $k_{\pm}=k\pm P/2$; $\Gamma_{\nu}^{V}$ is the Bethe-Salpeter amplitude for an on-shell $V$ meson – see, e.g., Ref. Xu et al. (2021); $S_{\mathpzc q}$ is the dressed $q$ quark propagator; – see, e.g., Ref. (Roberts et al., 2021, Sect. 2.C); and $\Gamma_{\mu}^{\gamma}$ is the associated dressed photon-quark vertex. Since we are focusing on photoproduction, then, with complete generality, one may use the Ball-Chiu form for $\Gamma_{\mu}^{\gamma}$ Ball and Chiu (1980).
The remaining element in Eq. (4) is $\Gamma^{\mathbb{P}}_{\alpha}=\gamma_{\alpha}\beta_{\mathpzc q}$, the Pomeron-quark vertex. Here, $\beta_{\mathpzc q}$ is a constant strength factor and $\gamma_{\alpha}$ is a standard Dirac matrix. We have verified that replacing $\gamma_{\alpha}$ by some more sophisticated structure, such as the Ball-Chiu vertex Ball and Chiu (1980), for instance, has no material influence: any impact can be absorbed into a modest modification of $\beta_{\mathpzc q}$.
Working in the center of three-momentum frame of the vector-meson–proton system, the relevant four-momenta are:
$q^{\mu}$ $=(\vec{q},i|\vec{q}|),$ (5a)
$P^{\mu}$ $=(\vec{P},i\sqrt{|\vec{P}|^{2}+m_{V}^{2}}),$ (5b)
$p_{1}^{\mu}$ $=(-\vec{q},i\sqrt{|\vec{q}|^{2}+m_{p}^{2}}),$ (5c)
$p^{\mu}_{2}$ $=(-\vec{P},i\sqrt{|\vec{P}|^{2}+m_{p}^{2}})\,$, (5d)
where the metric has uniform positive signature; $\vec{q}$ is the three-momentum of the on-shell photon; $\vec{P}$ is the vector-meson three-momentum; and
$|t|_{\rm min}=2|\vec{q}|(|\vec{P}|^{2}+m_{V}^{2})^{1/2}-m_{V}^{2}-2|\vec{q}||\vec{P}|\,.$ (6)
The differential cross section for exclusive vector meson photoproduction can now be expressed:
$\frac{d\sigma}{d\Omega}=\frac{1}{4\pi^{2}}\frac{m_{p}}{4W}\frac{|\vec{P}|}{k_{W}}\sum_{\text{proton spin}}(|{\cal I}_{1}|^{2}+|{\cal I}_{2}|^{2})\,$, (7)
where
$k_{W}=(W^{2}-m_{p}^{2})/(2m_{p})$ (8)
and $d\Omega$ is the differential solid-angle element defined by $\angle\vec{q}\vec{P}$. The corresponding differential cross section with respect to the momentum transfer is given by
$\frac{d\sigma}{d(-t)}=\frac{\pi}{|\vec{q}||\vec{P}|}\frac{d\sigma}{d\Omega}\,.$ (9)
The longitudinal cross section vanishes in photoproduction.
From the expressions sketched above, it is evident that the reaction model, combined with the Pomeron-exchange framework, offers a straightforward approach to studying vector meson photoproduction. The method is computationally economical and has been successfully applied to a broad class of processes; see, e.g., Ref. Donnachie and Landshoff (1992).
2.2 Pomeron in the reaction model
The reaction model involves two Pomeron elements. First, the Pomeron-proton content of the coupling is Pichowsky and Lee (1997); Donnachie and Landshoff (1984, 1987):
$\tilde{G}_{\alpha}(w^{2},t)=\gamma_{\alpha}F_{0}(t)3\beta_{\ell}G_{\mathbb{P}}(w^{2},t)\,$, (10)
where “33” enumerates the number of light valence degrees-of-freedom in the proton; $\beta_{\ell}$ is the strength of the associated Pomeron + light-quark coupling; and $F_{0}(t)$ is the empirical nucleon isoscalar elastic electromagnetic form factor. This last factor is well parametrised as follows on the spacelike low-$|t|$ domain relevant herein:
$F_{0}(t)=\frac{1-2.8t/[4m_{p}^{2}]}{1-t/[4m_{p}^{2}]}\frac{1}{(1-t/t_{0})^{2}}\,$, (11)
where $t_{0}=0.71\,$GeV2.
The second element in Eq. (10) is typically called the Pomeron propagator:
$G_{\mathbb{P}}(s,t)=\left[\frac{s}{s_{0}}\right]^{\alpha_{\mathbb{P}}(t)-1}\exp\left[-i\tfrac{\pi}{2}(\alpha_{\mathbb{P}}(t)-1)\right]\,$, (12)
with $\alpha_{\mathbb{P}}(t)=\alpha_{0}+\alpha_{1}t$, $s_{0}=1/\alpha_{1}$, and $\alpha_{0,1}$ fitted to selected small-$|t|$, large-$W$ meson photoproduction data within the context of Regge phenomenology: $\alpha_{0}$ is fixed by the energy dependence of the total cross section and $\alpha_{1}$ by the $t$-slope of the differential cross section – see, e.g., Refs. Collins (1977); Irving and Worden (1977). These parameters depend on whether the final-state vector meson is light or heavy Lee et al. (2022).
Table 1: Quark current-masses and parameters characterising the CSM interaction kernels for the mesons considered herein. The quark current masses are quoted at a renormalisation scale $\zeta=\zeta_{19}=19\,$GeV, i.e., in the far ultraviolet so as to avoid truncation ambiguity. The $u,d$ values were chosen to reproduce $m_{\pi}$, $f_{\pi}$; the kaon mass was used to fix $m_{s}$; $m_{c}$ was chosen to yield the $\eta_{c}$ mass; and the $\eta_{b}$ mass was used to fix $m_{b}$ Qin et al. (2018). Using one-loop evolution, the listed light quark masses correspond to $\zeta=\zeta_{2}=2\,$GeV values of $m_{\ell}^{\zeta_{2}}=0.0048\,$GeV, $m_{s}^{\zeta_{2}}=0.117\,$GeV; and regarding the heavy quarks, the listed values equate to the following Euclidean constituent quark masses Hecht et al. (2001), $M_{c}=1.41\,$GeV, $M_{b}=4.34\,$GeV. All four masses correspond to those typically quoted in connection with the given flavour and are a fair match with contemporary inferences Navas et al. (2024). The same Pomeron trajectory is used for both $\rho^{0}$ and $\phi$ photoproduction Pichowsky and Lee (1997), and another for $J/\psi$ and $\Upsilon$ (Chekanov et al., 2002, ZEUS). Pomeron-quark couplings, $\beta_{\mathpzc q}$, for both reaction models considered herein – Sects. 2.1, 2.3 – are determined as described in Sect. 2.4.
Meson Gluon model $\mathbb{P}$ trajectory $\mathbb{P}$-dyn $\mathbb{P}$-am
$m_{\zeta_{19}}$ [GeV] $D\omega$ $\text{GeV}^{3}$] $\omega$ [GeV] $N_{f}$ $\Lambda_{\text{QCD}}$ [GeV] $\alpha_{0}$ $\alpha_{1}$ $\text{GeV}^{-2}$] $\beta_{\mathpzc q}$ $\text{GeV}^{-1}$] $\beta_{\mathpzc q}$ $\text{GeV}^{-1}$]
$\rho^{0}_{\ell}$ 0.0034 $0.8^{3}$ 0.5 4 0.234 1.1 0.33 3.5 1.96
$\phi_{s}$ 0.083 $0.8^{3}$ 0.5 4 0.234 1.1 0.33 1.89 1.51
$J/\psi_{c}$ 0.89 $0.6^{3}$ 0.8 4 0.234 1.2 0.115 0.11 0.428
$\Upsilon_{b}$ 3.59 $0.6^{3}$ 0.8 5 0.36 1.2 0.115 0.016 0.638
2.3 VMD-like treatment of photon to meson transition
In the following, we will compare our predictions with those obtained using a simplified framework that replaces the calculated $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element by a phenomenologically constrained model. Sketched in Fig. 2 and identified as the $\mathbb{P}\,$-am model in Ref. Tang et al. (2024), this approach is obtained by beginning with the Ref. Sakinah et al. (2024) model; setting $v_{cN}=0$, so the quark-loop integration, based on a constituent quark model wavefunction, and VNV-N final state interactions (FSIs) do not contribute to the amplitude; and updating the Pomeron trajectories and Pomeron-quark couplings.
Figure 2: $\mathbb{P}$-am reaction model for $\gamma+p\to V+p$. Compared with our dynamical model, $\mathbb{P}$-dyn, sketched in Fig. 1, $\mathbb{P}$-am replaces the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element by a momentum-independent $\gamma\to V$ transition coupling (akin to vector meson dominance) along with a form factor, $F_{V}$, that expresses a Pomeron–vector-meson form factor.
The associated matrix element for diffractive vector meson photoproduction is
${\cal I}_{\mu}(W,t)\propto$ $\,\bar{u}(p_{2})\gamma_{\alpha}u(p_{1})\epsilon^{\lambda}_{\nu}(P)\Gamma_{\alpha,\mu\nu}G_{\mathbb{P}}(w^{2},t)$
$\times[\beta_{\mathpzc q}F_{V}(t)][\beta_{l}F_{0}(t)]\,$, (13)
with the tensor structure of the $\mathbb{P}$-$V$ vertex being Titov et al. (1998)
$\Gamma_{\alpha,\mu\nu}=q_{\alpha}\delta_{\mu\nu}-q_{\mu}\delta_{\alpha\nu}\,.$ (14)
The constants $\beta_{{\mathpzc q},\ell}$ fix the strengths of $\mathbb{P}$ interactions with the quarks in the vector mesons and the proton, respectively. Drawn from Ref. Sakinah et al. (2024), the phenomenological form factor
$F_{V}(t)=\frac{1}{m_{V}^{2}-t}\left(\frac{2\mu_{0}^{2}}{2\mu_{0}^{2}+m_{V}^{2}-t}\right),$ (15)
$\mu_{0}=1.1\,$GeV, replaces the momentum dependence of the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ loop and vastly simplifies its tensor structure.
2.4 Pomeron trajectories and couplings
In the subsequent calculations, we adopt Pomeron trajectories that are either well-established or reliably inferred, enabling accurate descriptions of existing experimental data Pichowsky and Lee (1997); Chekanov et al. (2002); Capua (2012). A common trajectory is used for light mesons and another for the heavy systems; see Table 1. The coupling $\beta_{u}$ is determined via a least-squares fit to the differential cross section for $\rho^{0}$ photoproduction (Derrick et al., 1995a, ZEUS1995). (The $\beta_{u}$ value used herein is 25% larger than that in Ref. Tang et al. (2024) as a consequence of correcting a coding error in that analysis.) The couplings $\beta_{s}$ and $\beta_{b}$ are obtained through fits to high-energy total cross-section data on $\phi$ Derrick et al. (1996) and $\Upsilon$ production Breitweg et al. (1998); Adloff et al. (2000); Chekanov et al. (2009); Aaij et al. (2015); Sirunyan et al. (2019); and the $\beta_{c}$ values are fixed, as in Ref. Tang et al. (2024), by requiring a best fit to the $W$-dependence of the $|t|\approx|t|_{\rm min}$ $J/\psi$ photoproduction differential cross section in Ref. (Chekanov et al., 2002, ZEUS 2002).
3 Photon to vector meson transition amplitude in CSMs
The photoproduction reaction mechanism sketched in Fig. 1 involves the dressed quark propagator, dressed photon-quark vertex; and vector-meson Bethe-Salpeter amplitude. Such quantities have long been the focus of CSM studies and many relevant reviews are available; see, e.g., Refs. Roberts (2015); Burkert and Roberts (2019); Bashir et al. (2012); Eichmann et al. (2016); Roberts et al. (2021); Binosi (2022); Ding et al. (2023); Ferreira and Papavassiliou (2023); Raya et al. (2024); Achenbach et al. (2025). Hence, today, these quantities are readily computed, so it is possible to arrive at parameter-free predictions for the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix elements, $t_{\mu\alpha\nu}(q,P)$ in Eq. (4). The key to such calculations is the quark+antiquark scattering kernel. For this, the leading-order (rainbow-ladder, RL) truncation Munczek (1995); Bender et al. (1996) is obtained by writing Maris and Roberts (1997):
$\mathscr{K}_{tu}^{rs}(k)$ $=\tilde{\mathpzc G}(y)[i\gamma_{\mu}\frac{\lambda^{a}}{2}]_{ts}[i\gamma_{\nu}\frac{\lambda^{a}}{2}]_{ru}T_{\mu\nu}(k)\,$, (16)
$k^{2}T_{\mu\nu}(k)=k^{2}\delta_{\mu\nu}-k_{\mu}k_{\nu}$, $y=k^{2}$. The tensor structure corresponds to Landau gauge, used because it is both a fixed point of the renormalisation group and that gauge for which corrections to RL truncation are least noticeable Bashir et al. (2009). In Eq. (3), $r,s,t,u$ represent colour, spinor, flavour matrix indices (as necessary).
A realistic form of $\tilde{\mathpzc G}(y)$ is explained in Refs. Qin et al. (2011); Binosi et al. (2015):
$\tilde{\mathpzc G}(y)$ $=\frac{8\pi^{2}}{\omega^{4}}De^{-y/\omega^{2}}+\frac{8\pi^{2}\gamma_{m}\mathcal{F}(y)}{\ln\big[\tau+(1+y/\Lambda_{\rm QCD}^{2})^{2}\big]}\,$, (17)
[…]
Applications to numerous systems and reactions have established Ding et al. (2023) that interactions in the class typified by Eqs. (3), (17) can unify the properties of many systems. Notably, when $\omega D=:\varsigma^{3}$ is held fixed, results for observable quantities remain practically unchanged under $\omega\to(1\pm 0.2)\omega$ Qin and Roberts (2020). Thus, the interaction in Eq. (17) is determined by just one parameter: $\varsigma^{3}:=\omega D$.
Table 2: Pion and vector meson masses and leptonic decay constants computed herein, using CSMs, compared with experimental values (Navas et al., 2024, PDG), quoted at a sensible level of precision. Regarding vector mesons, the mean absolute relative difference between prediction and experiment is 5(3)5(3)%.
Meson Mass [GeV] Decay constant [GeV]
Herein Expt. Herein Expt.
$\pi$ 0.138 0.138 0.092 0.092
$\rho^{0}$ 0.741 0.775 0.149 0.153
$\phi$ 1.086 1.019 0.183 0.168
$J/\psi$ 3.123 3.097 0.278 0.294
$\Upsilon(1s)$ 9.52 9.46 0.558 0.51
In the light-quark sector, $f=u,d,s$, $\varsigma_{f}=0.8\,{\rm GeV}$ delivers a good description of a range of pseudoscalar and vector meson static properties; see Table 2 and Fig. 3. These results were obtained as explained, e.g., in Refs. Chen et al. (2018); Ding et al. (2019); Xu et al. (2019). As noted above, the procedure involves solving sets of coupled gap and Bethe-Salpeter equations.
For heavier quarks, it is known that corrections to RL truncation are less important – see, e.g., Ref. Bhagwat et al. (2004) for illustration, so the interaction parameter should be closer to that used in more sophisticated kernels. This is discussed further in, e.g., Ref. (Xu et al., 2019, Sect. IIB), and explains the $c,b$ quark values listed in Table 1, which deliver the results listed in Table 2 and drawn in Fig. 3.
A
B
Figure 3: Pion and vector meson properties listed in Table 2, computed using CSMs with interaction parameters in Table 1 Qin et al. (2018); Xu et al. (2019); Yao et al. (2021): (A) Masses and (B) leptonic decay constants. Blue squares indicate CSM results; red triangles indicate experimental values at a sensible level of precision Navas et al. (2024).
4 Near-threshold slope parameter / meson-proton scattering lengths
[…]
$\frac{d\sigma^{\gamma p\to Vp}}{d\Omega}\bigg|_{\text{th}}$ $=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{1}{64\pi}|T^{\gamma p\to Vp}|^{2},$ (18)
$=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{\pi\alpha}{g_{V}^{2}}\frac{d\sigma^{Vp\to Vp}}{d\Omega}\bigg|_{\text{th}}$
$=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{\pi\alpha}{g_{V}^{2}}|\alpha_{Vp}|^{2},$
where $\alpha$ is the fine-structure constant, $\vec{q}$ and $\vec{P}$ are the initial and final momenta of the photon and vector meson, respectively – Eq.(5), and $\alpha_{Vp}$ is the model meson-proton scattering length. The VMD ingredient is the coupling $g_{V}$, which is defined by the vector meson’s electromagnetic decay width $\Gamma_{V\to e^{+}e^{-}}$:
$g_{V}^{2}=\frac{\pi\alpha^{2}m_{V}}{3\Gamma_{V\to\text{e}^{+}\text{e}^{-}}}.$ (19)
Our prediction for this differential cross section is given in Eq. (7). That is the formula we use to evaluate all results. However, by equating our predictions with Eq. (18) one can arrive at model-specific VMD values for the meson-proton scattering length. These can be used as single parameters with which to evaluate the veracity of existing such model dependent inferences of meson-proton scattering lengths Strakovsky et al. (2020a, b, 2021, 2025). Furthermore, obtained in this way, one need not interpret $|\alpha_{Vp}|$ as the VpV-p scattering length. Instead, it can be viewed as a near-threshold slope parameter. Then, the value has objective utility as a comparison between theory and experiment.
5 Results: cross sections
With the quark current-masses and interaction parameters, Table 1 – and the propagators and Bethe-Salpeter amplitudes they produce – validated in comparisons with experimental values for meson masses and decay constants, Table 2 and Fig. 3, we use these elements to calculate the matrix elements defined by Eqs. (3), (4).
Accounting for current-conservation and the fact that $P_{\nu}\Gamma_{\nu}^{V}=0$, then, in all cases, $t_{\mu\alpha\nu}^{V}$ involves $9$ independent scalar functions which modulate the strength of the allowed tensor structures:
$t_{\mu\nu\alpha}^{V}(q,P)=\sum_{i=1}^{9}F_{i}^{V}(q,P)\tau^{i}_{\mu\nu\alpha}(q,P).$ (20)
The nine tensors are listed explicitly in A.
Each scalar function in Eq. (20) is readily extracted using sensibly chosen projection operators. In practice, we find that the scalar function associated with the tensor structure
$\tau^{1}_{\mu\nu\alpha}(q,P)=q_{\alpha}(\delta_{\mu\nu}-P_{\mu}P_{\nu}/P^{2})$ (21)
gives the dominant contribution to the photoproduction cross sections in all cases: for heavy mesons, $J/\psi$, $\Upsilon$, all other tensors can be neglected; and for light mesons, $\rho$, $\phi$, it provides roughly 85% of the cross section, with the tensor
$\tau^{2}_{\mu\nu\alpha}(q,P)=P_{\alpha}(\delta_{\mu\nu}-P_{\mu}P_{\nu}/P^{2})$ (22)
contributing the remaining 15% – alone (5%) and via constructive interference with $\tau^{1}$ (10%). Notwithstanding these observations, all tensor structures are employed to produce the results reported herein.
For each vector meson, we calculate both the differential and total cross sections for exclusive photoproduction. The differential cross section depends on $W$, the centre-of-mass energy, and $t$, the square-momentum transfer. When relevant data are available, we examine both the $W$- and $t$-dependence. Often, the $W$ dependence is analysed at $|t|=|t|_{\text{min}}$, Eq. (6), which corresponds to forward scattering, i.e., the smallest possible value for which the vector meson can be produced leaving the proton in tact. The cross sections are largest on this domain and very sensitive to reaction details. For instance, within our approach, they are much affected by the quark-loop dynamics expressed by Fig. 1, as shown for the $J/\psi$ in Ref. Tang et al. (2024) and we will see again herein for other vector mesons. On the other hand, the $t$-dependence of the differential cross sections are usually mapped at fixed $W$. Near threshold, they, too, are sensitive to the vector meson production mechanism.
It is worth noting here that in fixed-target experiments, the photon energy is often used as the primary variable, viz. $E_{\gamma}(W)=k_{W}$ in Eq. (8).
5.1 $\rho$ photoproduction
Calculated using the Pomeron trajectory and quark–Pomeron couplings given in Table 1 – Row 1, our predictions for $\gamma+p\to\rho^{0}+p$ photoproduction are presented in Figs. 4 and 5.
Figure 4: $W$-dependence of total cross section for $\rho^{0}$ photoproduction, $W_{\rm th}^{\rho}=1.714\,$GeV. Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am. Data sources: black circles – (Struczinski et al., 1976, AHHM); cyan diamonds (fixed target) – (Aston et al., 1982, BCEGLMOPRS); green squares (Park et al., 1972; Ballam et al., 1972, SLAC); blue stars (Egloff et al., 1979, FNAL); red left and right triangles (Aid et al., 1996a; Andreev et al., 2020, H1); and red open squares – pseudodata inferred Strakovsky et al. (2025) from differential $\pi^{+}\pi^{-}p$ photoproduction cross sections measured with CLAS at JLab and analysed within the JM model Mokeev et al. (2016) (see text).
Figure 5: Differential cross section for $\rho^{0}$ photoproduction compared with data (Derrick et al., 1995a, ZEUS) (blue squares) in the range $60<W<$ $80$ GeV. Solid purple curve – $\mathbb{P}$-dyn; dashed black curve: $\mathbb{P}$-am. Both are computed at the average energy $\text{W}=70$ GeV. Pomeron trajectory and quark–Pomeron couplings used are given in Table 1 – Row 1.
In Fig. 4, we draw the total cross section on $W\in[W_{\rm th}^{\rho},500\,{\rm GeV}]$ along with data from a variety of sources (Struczinski et al., 1976, AHHM), (Aston et al., 1982, BCEGLMOPRS), (Park et al., 1972; Ballam et al., 1972, SLAC), (Egloff et al., 1979, FNAL), (Aid et al., 1996a; Andreev et al., 2020, H1), plus pseudodata obtained from differential $\pi^{+}\pi^{-}p$ photoproduction cross sections measured with the CLAS detector at JLab, analysed within the JLab-Moscow State University (JM) model Mokeev et al. (2016), and integrated over the differential solid angle. Both $\mathbb{P}$-dyn and $\mathbb{P}$- am reproduce the data on $W\gtrsim 10\,$GeV, which may therefore be identified as the domain of Pomeron dominance in this case. Furthermore, the $\mathbb{P}$-dyn model is satisfactory on $W\gtrsim 5\,{\rm GeV}$, highlighting the importance of a sound description of the quark loop in Fig. 1. However, both models are equally poor on $W\in[W_{\rm th}^{\rho},5\,{\rm GeV}]$. This is because the proton and $\rho$ meson share valence quarks; hence, many additional (nondiffractive) exchange contributions are possible in the photoproduction reaction. For instance, a leading-order calculation using a phenomenological meson-exchange model Sato and Lee (1996) is able to reproduce the low-$W$ data (Pichowsky and Lee, 1997, Fig. 15). Such contributions decay rapidly with $W$, so that the diffractive $\mathbb{P}$-dyn reaction model becomes valid on $W\gtrsim 5\,$GeV.
Figure 5 depicts the differential cross section. It compares $\mathbb{P}$-dyn and $\mathbb{P}$-am predictions with data from Ref. (Derrick et al., 1995a, ZEUS) in the range $60<\text{W}<$ $80$ GeV. Our results are evaluated at the average experimental energy, $W=70$ GeV. Recalling that the Pomeron trajectory parameters in Table 1 – Row 1 were taken from Ref. Pichowsky and Lee (1997) and only the quark-Pomeron couplings were tuned to the data in Fig. 4 – see Sec. 2.4 – then it becomes clear that the quark-loop dynamics in Fig. 1 has an impact on the differential cross section even far above threshold, $W\gg W_{\rm th}^{\rho}$.
Regarding $\omega$-meson photoproduction, owing to its comparable structure, similar features and explanations pertain to the cross sections Yu et al. (2019).
Figure 6: $W$-dependence of total cross section for $\phi$ photoproduction, $W_{\rm th}^{\phi}=1.958\,$GeV. Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am. Data: blue stars – (Egloff et al., 1979, FNAL); cyan triangles – (Busenitz et al., 1989, E401); black circles – (Derrick et al., 1996, ZEUS); and green open squares “pseudodata” Strakovsky et al. (2020b), obtained by interpolating CLAS differential cross sections (Dey et al., 2014, CLAS) as described in the text.
5.2 $\phi$ photoproduction
Using the Pomeron trajectory and quark-Pomeron couplings given in Table 1 – Row 2, we obtain the predictions for $\gamma+p\to\phi+p$ photoproduction cross sections presented in Figs. 6 – 10. An important difference between this reaction and $\rho^{0}$ photoproduction is that the $\phi$-meson contains no valence quarks in common with the proton; hence, fewer nondiffractive processes are possible.
The total cross section, $\sigma_{\gamma p\to\phi p}$, is drawn in Fig. 6. Comparing the $\mathbb{P}$-dyn and $\mathbb{P}$-am predictions with data, it is evident that the dynamical quark loop associated with Fig. 1 is crucial in order to obtain agreement with data. This is especially true near threshold. Notably, very near threshold, direct measurements of $\sigma_{\gamma p\to\phi p}$ are unavailable. To fill that gap, Ref. Strakovsky et al. (2020b) generated pseudodata by interpolating differential cross sections measured at JLab (Dey et al., 2014, CLAS) and integrating over the differential solid angle. The $\mathbb{P}$-dyn prediction aligns well with this pseudodata, too; although it misses the “bump” apparent in the pseudodata. We note that this bump may be merely an artefact of the pseudodata construction procedure.
Figure 7: $E_{\gamma}$ dependence of differential cross section for $\phi$ photoproduction, evaluated at $|t|=|t|_{\text{min}}$, Eq. (6), corresponding to forward-angle scattering. Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am. Data: black stars – (Ballam et al., 1973, SLAC); black up triangles – (Besch et al., 1974, Bonn); cyan squares – (Behrend et al., 1978, DESY); green – (Barber et al., 1982, Daresbury); black down triangles – (Barth et al., 2003, SAPHIR); blue open squares – (Mibe et al., 2005, LEPS); red circles – (Seraydaryan et al., 2014, CLAS).
The $E_{\gamma}$ dependence of the forward-angle $\phi$ photoproduction differential cross section is drawn in Fig. 7. Compared with data, the $\mathbb{P}$-dyn reaction model delivers a better global description than $\mathbb{P}$-am; hence, our dynamical treatment of the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to\phi$ transition is once more seen to be important.
A
B
Figure 8: $|t|$ dependence of differential cross section for $\phi$ photoproduction at photon energies: $E_{\gamma}=3.3\,$GeV ($W\approx 2.66$ GeV) and $E_{\gamma}=3.6\,$GeV ($W\approx 2.76$ GeV). Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$ - am. Data: red up-triangles – (Dey et al., 2014, CLAS 2014); blue squares – (Barber et al., 1982, Daresbury) (Panel A); red down-triangles – (Anciant et al., 2000, CLAS 2000) (Panel B).
The $|t|$ dependence of the $\phi$ photoproduction differential cross section is depicted in Fig. 8 for two different but low photon energies. On $|t|\lesssim 2\,$GeV2, the $\mathbb{P}$-dyn reaction model provides a good description of the data (Dey et al., 2014, CLAS 13), (Barber et al., 1982, Daresbury), (Anciant et al., 2000, CLAS 00). However, the reaction is not purely diffractive on $|t|\gtrsim 2\,$GeV2: on this domain, one sees structure, likely associated with baryon resonance excitation. The mismatch between $\mathbb{P}$-am and data on the entire domain depicted highlights the importance of a realistic description of the $\gamma\to\bar{q}q+\mathbb{P}\to V$ transition, Fig. 1.
Figure 9: Differential cross sections for $\phi$ photoproduction as a function of $|t|-|t|_{\text{min}}$, Eq. (6), at low photon energies, $E_{\gamma}$. Solid purple curve – $\mathbb{P}$-dyn. Data: red up-triangles – (Dey et al., 2014, CLAS 2014); black down-triangles – (Barth et al., 2003, SAPHIR).
Figure 10: Differential cross sections for $\phi$ photoproduction as a function of $|t|-|t|_{\text{min}}$ at low photon energies, $E_{\gamma}$. Solid purple curve – $\mathbb{P}$-dyn. Data: red up-triangles – (Dey et al., 2014, CLAS 2014); blue circles – (Mibe et al., 2005, LEPS).
Figures 9, 10 compare $\mathbb{P}$-dyn predictions for the $|t|$-dependence of the $\phi$ photoproduction differential cross sections with data available at lower photon energies than those in Fig. 8, viz. Refs. (Dey et al., 2014, CLAS 13), (Barth et al., 2003, SAPHIR), (Mibe et al., 2005, LEPS). Regarding Fig. 9, in this case, too, the agreement between our reaction model and data for near-forward scattering is good. The mismatch at larger $|t|$ was already discussed in connection with Fig. 8. Concerning Fig. 10, the $\mathbb{P}$-dyn provides a good description of the differential cross section except, perhaps, in the neighbourhood $E_{\gamma}\simeq 2\,$GeV, whereupon some nondiffractive processes may be contributing somewhat to the production process.
Figure 11: Differential cross-section for $\gamma p\to J/\psi p$ on $|t|/m_{p}\simeq 0$ (forward). Solid purple curve: $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am. (Pomeron parameters in Table 1 – Row 3.) Data from Ref. (Chekanov et al., 2002, ZEUS 2002).
5.3 $J/\psi$ photoproduction
The reaction model characterised by the sketch in Fig. 1, $\mathbb{P}$-dyn, was applied to $J/\psi$ photoproduction in Ref. Tang et al. (2024). Herein, for completeness, we recapitulate and somewhat extend the results described in that study. Of course, the $J/\psi$ contains no valence degrees of freedom in common with the proton; so, it has long been widely held Krein et al. (2018); Lee et al. (2022) that the dominant mechanisms underlying $J/\psi$ photoproduction must involve some manifestations of gluon physics within the target proton and, perhaps, the incipient vector meson, and/or the exchange of (perhaps infinitely many, correlated) gluons between the partonic constituents of each. (This latter is a Pomeron-like effect.)
The $J/\psi$ process is topical because some have argued Kharzeev (1996) that there is a connection between near-threshold $J/\psi$ photoproduction and the in-proton expectation value of the QCD trace anomaly; hence, insights into the character of emergent hadron mass (EHM) Roberts (2017); Krein and Peixoto (2020); Roberts et al. (2021); Binosi (2022); Ding et al. (2023); Ferreira and Papavassiliou (2023); Salmè (2022); Carman et al. (2023); Achenbach et al. (2025). This possibility has served as a motivation for new high-energy, high-luminosity accelerator facilities Chen et al. (2020); Anderle et al. (2021); Abdul Khalek et al. (2022). However, any trace-anomaly connection now appears remote Du et al. (2020); Xu et al. (2021); Sun et al. (2022); Sakinah et al. (2024). Notwithstanding these remarks, there is a continuing desire to interpret $\gamma+p\to V+p$ photoproduction data in terms of gluon physics; hence, an elucidation of the underlying reaction mechanism is crucial.
As discussed in Sect. 2.4 and in Ref. Tang et al. (2024), the Pomeron–$c$-quark coupling was fixed by requiring a best least-squares fit to the $W$-dependence of the $|t|\approx|t|_{\rm min}$ $J/\psi$ photoproduction differential cross section in Ref. (Chekanov et al., 2002, ZEUS 2002). The results are displayed in Fig. 11. From this point, all $J/\psi$ results are predictions of our reaction model. Thus, it is significant that, as illustrated inFig. 12, our reaction model provides a good description of the $(W,t)$ dependence of the differential cross sections reported in Ref. (Adloff et al., 2000, H1 2000). Evidently, with increasing $|t|$ high-$W$ data retain some sensitivity to the dynamical quark loop in Fig. 1: compare $\mathbb{P}$-dyn with $\mathbb{P}$-am on $|t|\gtrsim 1\,$GeV2.
Figure 12: Predicted $|t|$ dependence of differential cross section for $J/\psi$ photoproduction at selected $W$ values well above threshold. Solid purple curve – $\mathbb{P}$-dyn, W=140W=140\,GeV; dashed purple curve – $\mathbb{P}$-dyn, W=40W=40\,GeV; dot-dashed black curve – $\mathbb{P}$- am, W=140W=140\,GeV; dotted black curve – $\mathbb{P}$- am, W=40W=40\,GeV. Data, with $\bar{W}$ being experimental mean $W$: blue up-triangles – (Adloff et al., 2000, H1 2000); red down-triangles – (Adloff et al., 2000, H1 2000); cyan circles – (Derrick et al., 1995b, ZEUS 1995).
Figure 13: Differential cross-section for $\gamma p\to J/\psi p$. Solid purple curve – $\mathbb{P}$-dyn; dotted purple curve – $\mathbb{P}$-am; dot-dashed blue curve – GPD model Guo et al. (2023); dashed green curve – Pomeron + $J/\psi N$ scattering model Sakinah et al. (2024). (Predictions herein obtained with Pomeron parameters in Table 1 – Row 3.) Data: $4.43<W/{\rm GeV}<4.8$(Ali et al., 2019, GlueX  2019); $4.38<W/{\rm GeV}<4.63$(Adhikari et al., 2023, GlueX 2023); $4.41<W/{\rm GeV}<4.54$(Duran et al., 2023, $J/\psi$-007). (All theory curves computed using W=4.5W=4.5\,GeV.)
Figure 13 depicts our predictions for the $\gamma p\to J/\psi p$ differential cross section as a function of $|t|$ at energies not too far above threshold, $W_{\rm th}^{J/\psi}=4.04\,$GeV. Evidently, on this $W$ domain, the momentum-dependence of the $\gamma\to c\bar{c}+\mathbb{P}\to J/\psi$ transition loop in Fig. 1 has a significant influence. Again, this is highlighted by the difference between the $\mathbb{P}$-dyn and $\mathbb{P}\,$-am predictions.
Along with our parameter-free predictions, we also draw the result obtained using a GPD model Guo et al. (2023), which involves three parameters and assumes specific, simple forms for the two gluon gravitational form factors: the parameters were fixed via a least-squares fit to the data in Refs. Adhikari et al. (2023); Duran et al. (2023). Citing Ref. Tang et al. (2024), quantitatively, the GPD model’s description of $J/\psi$-007 data Duran et al. (2023) is better than its match with the GlueX points Adhikari et al. (2023). In this outcome, there is a signal of overfitting, the probability of which is increased by the result for $d\sigma/dt$ being a concave function on the entire plotted $|t|$ domain, with a global maximum on $|t|\simeq 0.7\,$GeV2. This is qualitatively inconsistent with the other reaction models, which deliver a differential cross-section that is a monotonically decreasing convex function.
The remaining curve in Fig. 13 is the result from Ref. Sakinah et al. (2024), which argues that $J/\psi$ photoproduction should be described by Pomeron exchange augmented by a $c$-quark+nucleon potential, $v_{cN}$, that drives $J/\psi\,N$ scattering and FSIs, with the FSIs dominating near threshold. This result is similar to our favoured $\mathbb{P}$-dyn prediction on $|t|\lesssim 2.5\,$GeV2 and thereafter becomes harder. Again citing Ref. Tang et al. (2024), quantitatively, the FSIs in the $\mathbb{P}+v_{cN}$ reaction model improve the description of GlueX data but degrade that of $J/\psi$-007. Looking harder, one finds that the most reasonable description of available near-threshold $d\sigma/dt$ data is provided by our $\mathbb{P}$-dyn reaction model: this parameter-free prediction delivers a monotonically decreasing convex differential cross-section and a $\chi^{2}/$degree-of-freedom that does not vary excessively across the data sets Tang et al. (2024).
These observations show that whilst modern Poincaré covariant calculations do indicate that the proton mass radius, $r_{m}$, is smaller than its charge radius Yao et al. (2025), owing to the different character of the effective probe Xu et al. (2023), one is not justified in drawing any link between the Ref. (Duran et al., 2023, $J/\psi$-007) data and a measurement of $r_{m}$ or, indeed, any other quantity inferred via a GPD-specific interpretation of those or other data.
A
B
C
Figure 14: $|t|$ dependence of $J/\psi$ photoproduction differential cross sections near threshold, $W_{\rm th}^{J/\psi}=4.04\,$GeV. Solid purple curve – $\mathbb{P}$-dyn; dotted purple – $\mathbb{P}$-am. The different panels depict, respectively photon energies $E_{\gamma}=8.93,9.85,10.82\,$GeV, which correspond to W=4.2,4.4,4.6W=4.2,4.4,4.6\,GeV. Data: black circles – (Adhikari et al., 2023, GlueX); cyan squares and triangles – (Duran et al., 2023, $J/\psi$-007).
We take a closer look at near-threshold $J/\psi$ photoproduction in Fig. 14 by exploring the $(E_{\gamma},|t|)$ dependence of the differential cross-section. Very close to threshold, Fig. 14 A, the Ref. (Adhikari et al., 2023, GlueX) data suggest a role for some nondiffractive contributions to the reaction on $|t|\gtrsim 3\,$GeV2. This domain is not reached by the Ref. (Duran et al., 2023, $J/\psi$-007) data: on their domain of coverage, those data are entirely consistent with our $\mathbb{P}$-dyn reaction model. Reviewing Figs. 14 B, C, it is evident that, as one would expect, the importance of any nondiffractive contributions to the differential cross section diminishes with increasing $E_{\gamma}$.
A
B
Figure 15: $W$ dependence of total $J/\psi$ photoproduction cross section, $W_{\rm th}^{J/\psi}=4.04\,$GeV. Panel A. Near-threshold region, $W_{\rm th}^{J/\psi}<W<5.5\,$GeV. Panel B. Entire coverage of available data, $W_{\rm th}^{J/\psi}<W<2\,$TeV. Solid purple curve – $\mathbb{P}$-dyn; dashed black – $\mathbb{P}$-am. Data: blue stars – (Adloff et al., 2000, H1 2000); cyan up-triangles – (Chekanov et al., 2002, ZEUS 2̇002); red right-triangles – (Aaij et al., 2014, LHCb 2014); red circles – (Adhikari et al., 2023, GlueX 2023); green squares – (Aaij et al., 2025, LHCb 2024); black diamonds – pre-2000 Camerini et al. (1975); Gittelman et al. (1975); Shambroom et al. (1982); Derrick et al. (1995b); Frabetti et al. (1993); Aid et al. (1996b); Amarian (1999).
Our predictions for the $W$ dependence of the total $\gamma p\to J/\psi p$ photoproduction cross section are displayed in Fig. 15. As explained in Ref. Tang et al. (2024), the $\mathbb{P}$-dyn reaction model provides a good description of all available data (Adloff et al., 2000, H1 2000), (Chekanov et al., 2002, ZEUS 2̇002), (Aaij et al., 2014, LHCb 2014), (Adhikari et al., 2023, GlueX 2023), (Aaij et al., 2025, LHCb 2024), and pre-2000 Camerini et al. (1975); Gittelman et al. (1975); Shambroom et al. (1982); Derrick et al. (1995b); Frabetti et al. (1993); Aid et al. (1996b); Amarian (1999), which today covers a domain from threshold to $W\simeq 2\,$TeV. Near threshold, a realistic description of the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element in Fig. 1 is crucial in achieving agreement with data: the $\mathbb{P}$-am model fails on this domain, becoming viable only on $W\gtrsim 5\,$GeV.
Emulating Ref. (Tang et al., 2024, Fig. 6), Fig. 15 also includes comparisons with other reaction models. Regarding the GPD model, the best-fit curve from Ref. Guo et al. (2023) is drawn: the framework is inapplicable on $W\gtrsim 5\,$GeV; hence, it cannot meet a requirement for providing a uniformly applicable approach. Furthermore, even on the subdomain of assumed applicability, the description of (Ali et al., 2019; Adhikari et al., 2023, GlueX) data does not match that produced by the $\mathbb{P}$-dyn reaction model, viz. the approach advocated herein. These facts further impair attempts to connect near threshold $J/\psi$ photoproduction data with the in-proton gluon GPD.
A
B
Figure 16: $|t|$ dependence of differential cross section for $\Upsilon$ photoproduction. Panel A. W=11W=11\,GeV. Panel B. $\mathbb{P}$-dyn: solid purple – $W=100\,{\rm GeV}$; long-dashed purple – 500500\,GeV; dot-dashed purple – 11\,TeV. $\mathbb{P}$-am: dashed black – $W=100\,{\rm GeV}$; dot-dot-dashed black – 500500\,GeV; dotted black – 11\,TeV.
On the other hand, the reaction model in Ref. Sakinah et al. (2024), delivers a good description of available total cross-section data on the domain $W_{\rm th}^{J/\psi}<W<300\,$MeV. In this case, considering also the model’s results for the differential cross section – see Fig. 12, it is the link drawn between such agreement and dominance of FSIs near threshold that is debatable. The connection relies on a particular choice for the Pomeron trajectory. If one instead uses the trajectory we have described – Table 1 - Row 3 – namely, that associated with $\mathbb{P}$-dyn, then an equivalent description of data is possible without calling upon FSIs.
5.4 $\Upsilon$ photoproduction
To complete the picture, it remains only to consider $\Upsilon$ photoproduction, the threshold for which is $W_{\rm th}^{\Upsilon}=10.40\,$GeV. Today, there are no differential cross section data and total cross sections are only available on $100\,\lesssim W/{\rm GeV}\lesssim 2\,000$. Owing to the large mass of the valence $b$, $\bar{b}$ quarks, there are attempts to employ perturbative QCD concepts as the basis for explaining available data; see, e.g., Refs. Eskola et al. (2023); Penttala and Royon (2025); Xie and Gonçalves (2025). As we shall see, this is by no means mandatory.
Our predictions for the $\gamma p\to\Upsilon p$ differential cross section are drawn in Fig. 16. Owing to the large $\Upsilon$ mass, $t_{\rm min}\approx 2.2\,$GeV2 near-threshold. Naturally, it diminishes with increasing $W$. Regarding Fig.  16 A, it is apparent that even for $\Upsilon$ photoproduction, the dynamical quark loop in Fig. 1 plays a material role near threshold. However, as in the other cases considered herein, precise data would be necessary to distinguish between competing reaction mechanisms on this domain. Turning to Fig. 16 B, it is evident that, with increasing $W$, the Pomeron trajectory dominates the differential cross section on the diffractive domain. It is worth stressing here that we use the same Pomeron trajectory for both $J/\psi$ and $\Upsilon$ production.
Figure 17: $W$ dependence of total cross section for $\Upsilon$ photoproduction, with $W_{\rm th}^{\Upsilon}=10.40\,$GeV. Solid purple curves – $\mathbb{P}$-dyn; dashed black – $\mathbb{P}$-am. Data: cyan squares – (Breitweg et al., 1998, ZEUS 1998); green up-triangles – (Adloff et al., 2000, H1 2000); black down-triangles – (Chekanov et al., 2009, ZEUS 2009); blue diamonds – (Aaij et al., 2015, LHCb 2015); red stars – (Sirunyan et al., 2019, CMS 2018).
The total cross section for $\Upsilon$ photoproduction is displayed in Fig. 17. The $\mathbb{P}$-dyn and $\mathbb{P}$-am reaction models yield equivalent results; so one may conclude that on the domain covered by extant experiments, simple Pomeron exchange is sufficient to describe the data. Precise data closer to threshold would be valuable. This may be expected from experiments at future electron ion colliders Chen et al. (2020); Anderle et al. (2021); Abdul Khalek et al. (2022).
Figure 18: Total cross sections for all vector meson photoproduction reactions considered herein. Each curve is the associated $\mathbb{P}$-dyn prediction. The data sources are recorded in Figs. 4, 6, 12, 17.
[…]
A Eq. (23a) Eq. (23b) Expt.
$V$ XX/nb $\epsilon$ $\sigma_{P}$/nb $\epsilon$ $\epsilon$
$\rho^{0}$ $4076$ 0.27 5611 0.21 $0.17\pm 0.01^{+0.04}_{-0.03}$
$\phi$ $234.2$ 0.26 430.0 0.20 $\approx\epsilon_{\rho}$ Capua (2012); Lukashin et al. (2001)
$J/\psi$ $2.136$ 0.73 2.178 0.73 $0.67\pm 0.03$
$\Upsilon$ $0.0064$ 0.73 0.0066 0.73 $0.77\pm 0.14$
B Eq. (23a) Eq. (23b) Expt.
$V$ XX/nb $\epsilon$ $\sigma_{P}$/nb $\epsilon$ $\epsilon$
$\rho^{0}$ $3368$ 0.32 4151 0.27 $0.17\pm 0.01^{+0.04}_{-0.03}$
$\phi$ $268.6$ 0.30 343.7 0.25 $\approx\epsilon_{\rho}$ Capua (2012); Lukashin et al. (2001)
$J/\psi$ $2.099$ 0.74 2.139 0.74 $0.67\pm 0.03$
$\Upsilon$ $0.0067$ 0.74 0.0066 0.73 $0.77\pm 0.14$
6 Results: power-law behavior of vector meson photoproduction cross sections
Figure 18 records the total cross sections for all vector meson photoproduction reactions considered herein, depicting $\mathbb{P}$-dyn predictions vs. data. In the analysis of data, it is common to report power-law exponents that characterise the large-$W$ behaviour of such cross sections. They are typically extracted using one of the following formulae Donnachie and Landshoff (1992); Klein et al. (2017):
$\sigma(W)$ $=XW^{\epsilon},$ (23a)
$\sigma(W)$ $=\sigma_{P}\left[1-\frac{(m_{p}+m_{V})^{2}}{W^{2}}\right]^{2}W^{\epsilon}.$ (23b)
NB. Equation (23b) is applicable $\forall W$, whereas Eq. (23a) can only be used on some domain well above threshold, viz. $W>W_{\rm min}\gg W_{\rm th}$.
For the $\mathbb{P}$-dyn reaction model, values of the fitting parameters are listed in Table 3 A and compared with empirical inferences. In all cases, the comparison is good when using Eq. (23b); and Eq. (23a) serves well for heavy mesons.
Employing analogous $\mathbb{P}$-am results from Figs. 4, 6, 14, 16, one obtains the exponents listed in Table 3 B. In this case, a favourable comparison with empirical inferences is only achieved for heavy mesons, $V=J/\psi,\Upsilon$.
Table 4: VMD-based scattering lengths – Eq. (18) – inferred from near-threshold $V$-meson photoproduction differential cross sections calculated using the $\mathbb{P}$-dyn and $\mathbb{P}$-am reaction models. $q_{\rm th}$ is the photon momentum at threshold in the centre-of-mass frame. The value of $|\alpha_{\Upsilon p}^{\rm VMD}|$ was estimated in Ref. Strakovsky et al. (2021) by using near-threshold pseudodata created using the GPD model in Ref. Guo et al. (2021), which is an earlier version of the GPD model discussed above Guo et al. (2023): the GPD reaction model delivers results that disagree with ours by an order of magnitude and more.
$V$ $m_{V}$/GeV $q_{th}$/GeV $\Gamma_{V\to e^{+}e^{-}}$/keV $|\alpha_{Vp}^{\rm dyn}|$/fm $|\alpha_{Vp}^{\rm am}|$/fm $|\alpha_{Vp}^{\rm VMD}|$/fm
$\rho^{0}\$0.775 0.6 6.97 0.136 0.072 $0.23\pm 0.03$ Strakovsky et al. (2025)
$\phi^{0}\$1.019 0.75 1.27 0.087 0.048 $0.063\pm 0.010$ Strakovsky et al. (2020b)
$J/\psi\$3.097 1.91 5.53 $3.08\times 10^{-3}$ $1.88\times 10^{-3}\$$3.08(55)\times 10^{-3}$ Strakovsky et al. (2020a)
$\Upsilon\$9.46 5.16 1.34 $5.87\times 10^{-5}$ $1.86\times 10^{-5}\$$0.51(3)\times 10^{-3}$ Strakovsky et al. (2021)
In the standard Pomeron-alone framework, the exponent should be Donnachie and Landshoff (1992):
$\epsilon\approx 4[\alpha_{\mathbb{P}}(0)-1]\,.$ (24)
Consideration of Eqs. (3), (7), (12) indicates that a better estimate is provided by the following formula:
$\epsilon(t_{\rm min}(W))\approx 4[\alpha_{\mathbb{P}}(t_{\rm min}(W))-1]\,$, (25)
where $|t_{\rm min}(W)|$ decreases with increasing $W$. Notwithstanding Eq. (25), phenomenological analyses of data use Eqs. (23); so return a $W$-independent exponent, which must therefore be some mean-$\epsilon$, averaged over the measurement domain. Working from Table 1, Eq. (24) yields $\epsilon_{\rho}=\epsilon_{\phi}=0.4$ and $\epsilon_{J/\psi}=\epsilon_{\Upsilon}=0.8$. Evidently, considering Table 3, whilst Eq. (24) is a fair approximation for heavy meson data, $V=J/\psi,\Upsilon$, it delivers a poor estimate for light meson data, $V=\rho,\phi$.
We note that both the $\mathbb{P}$-dyn and $\mathbb{P}$-am reaction models deliver exponents in qualitative agreement with Eqs. (24), (25): since $\alpha_{0}^{\rho}=\alpha_{0}^{\phi}$, $\alpha_{0}^{J/\psi}=\alpha_{0}^{\Upsilon}$ in Table 1, and $|t_{\rm min}|$ is small, then $\epsilon_{\rho}\approx\epsilon_{\phi}$ and $\epsilon_{J/\psi}\approx\epsilon_{\Upsilon}$.
[…]
This discussion highlights that structure in the $\gamma\to V$ transition component of the $\gamma p\to Vp$ matrix element has an impact on the exponent value. Naturally, the effects are greater for the $\mathbb{P}$-dyn reaction model and, indeed, are just what is needed to deliver agreement with the empirically deduced exponent for light mesons. As we have shown above, such corrections diminish at the typically larger $W$ values associated with $V=J/\psi,\Upsilon$ photoproduction.
7 Results: near-threshold slope parameter / vector meson-proton scattering lengths
As noted in Sect. 4, despite its weaknesses Xu et al. (2021), VMD remains a widely used phenomenological tool; and within that framework, there is a simple relationship between the differential cross section for $\gamma p\to Vp$ and the VpV-p scattering length, namely, Eq. (18). Our prediction for the $\gamma p\to Vp$ differential cross section is given in Eq. (7) and this is the formula we used above to calculate all results.
Figure 19: VMD-based scattering lengths – Eq. (18) – inferred from near-threshold vector meson photoproduction differential cross sections. Pictorial representation of the results listed in Table 4.
Here, for the sake of a comparison between our predictions and the model-specific VMD-values of meson-proton scattering lengths inferred in Refs. Strakovsky et al. (2020a, b, 2021, 2025), we equate our near-threshold results with Eq. (18) to arrive at a crude estimate of $|\alpha_{Vp}|$. The results are listed in Table 4 and plotted in Fig. 19. The differences between $\mathbb{P}$-dyn and $\mathbb{P}$-am predictions again highlight the importance of the dynamical quark loop – Fig. 1 – near threshold. It is worth recording the following pattern in our predictions:
$|\alpha_{\Upsilon p}|\ll|\alpha_{J/\psi p}|\ll|\alpha_{\phi p}|<|\alpha_{\rho^{0}p}|\,.$ (26)
The $\mathbb{P}$-dyn results may be described by
$|\alpha_{Vp}|\approx{\mathpzc a}_{0}/m_{V}^{3},\quad{\mathpzc a}_{0}=(0.41\,{\rm GeV})^{3}\,{\rm fm}.$ (27)
As observed in Sect. 4, one need not interpret $|\alpha_{Vp}|$ as the VpV-p scattering length; instead, it can simply be viewed as a near-threshold slope parameter, in which case the value has objective utility as a comparison between theory and experiment.
Regarding $V=\rho^{0}$, there is qualitative agreement between our near-threshold slope and that associated with available data. One should not expect better because, as explained in Sect. 5.1, nondiffractive processes are important for $\rho$-meson photoproduction and the empirically inferred value is based on pseudodata Strakovsky et al. (2025). Turning to $\phi$ photoproduction, Sect. 5.2 describes good agreement between our predictions and data, so our value for the $\phi$ slope parameter should be reliable. Hence, the agreement with data is good, especially when one considers the scatter in available near threshold data; see Figs. 7 – 10. Owing to the heightened interest in $J/\psi$ photoproduction, a good deal of modern data is available near threshold. Our prediction for the associated slope parameter agrees well with the value deduced from the new data.
The outlier in Table 4 is the slope parameter associated with near threshold $\Upsilon$ photoproduction. In our view, given its success with $J/\psi$ photoproduction, the $\mathbb{P}$-dyn reaction model should be reliable for $\Upsilon$, too. However, its prediction is an order of magnitude and more smaller than that listed in Ref. Strakovsky et al. (2021). We judge that the explanation for this mismatch is straightforward. The estimate in Ref. Strakovsky et al. (2021) is based on pseudodata produced near threshold using the GPD model in Ref. Guo et al. (2021), which is an earlier version of the GPD model discussed above Guo et al. (2023). As observed in Sect. 5.3 and Ref. Tang et al. (2024), the GPD model provides a poor representation of near-threshold $J/\psi$ data; so one should not expect it to be reliable for $\Upsilon$ photoproduction. Data from future electron ion collider experiments may be expected to resolve this issue. Meanwhile, we present our prediction as a solidly benchmarked estimate.
8 Summary and perspective
We described and employed a reaction model, $\mathbb{P}$-dyn, for $\gamma+p\to V+p$ photoproduction, $V=\rho^{0},\phi,J/\psi,\Upsilon$, that exposes the quark-antiquark content of the dressed photon in making the transition $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$, where $\mathpzc q$} depends on $V$, and couples the intermediate ${\mathpzc q}\bar{\mathpzc q}$ system to the proton’s valence quarks via Pomeron exchange [Fig. 1]. The parameters in this model are those characterising the Pomeron trajectories for light (two) and heavy mesons (two) and the Pomeron-quark couplings (four). With those parameters fitted to high-$W$ data [Sect. 2.4], $\mathbb{P}$-dyn provides a uniformly good description of a large number of differential and total cross-sections for $V$-meson photoproduction from the proton on the entire kinematic ranges measured to date [Sect. 5]. In all cases, the quality of the description is as good as or better than that provided by alternative models. The quark loop in Fig. 1 plays a key role in the success of the $\mathbb{P}$-dyn reaction description.
There is one caveat, however. Namely, as one moves away from the forward limit, nondiffractive mechanisms can play a role in the photoproduction process. Since the $\rho^{0}$ contains valence degrees of freedom in common with the proton, that is most apparent for $\rho^{0}$ production. In this case, the $\mathbb{P}$-dyn reaction model is insufficient on a domain that extends from threshold, $W_{\rm th}^{\rho}$, to $W\approx 5\,$GeV, where $W$ is the centre of mass energy [Fig. 4].
Working with $\mathbb{P}$-dyn predictions, we extracted the power-law exponents that are often used empirically to characterise the large-$W$ behaviour of total cross sections for $\gamma+p\to V+p$ photoproduction [Sect. 6]. Our results agree well with phenomenological inferences from data. For light mesons, the exponents do not take the values suggested by simple Pomeron pictures. Again, the quark loop in Fig. 1 plays a key role in these outcomes.
We also used the $\mathbb{P}$-dyn predictions to extract near-threshold slope parameters that can be used to characterise the differential cross sections for $\gamma+p\to V+p$ photoproduction [Sect. 7]. Owing to the role of nondiffractive processes, we judge that the $\rho^{0}$ value is underestimated. On the other hand, the predictions for $V=\phi,J/\psi,\Upsilon$ are solidly benchmarked; hence, should serve as realistic estimates.
Acknowledging the $\rho$-meson caveat, our $\mathbb{P}$-dyn reaction model delivers a uniformly good, unified description of forward $V=\rho^{0},\phi,J/\psi,\Upsilon$ meson photoproduction from relevant thresholds to very high energies. A key to this success is a realistic nonperturbative description of the $\gamma\to V$ transition, which is especially important near threshold but also plays a role at higher energies. Apart from the interaction dynamics expressed in treating that loop, the Pomeron provides the only link to “gluon” physics in the reaction model and, so far is now known, that mechanism is unrelated to in-proton gluon distributions. Hence, at present, any attempt to constrain such distributions using near-threshold $\gamma+p\to V+p$ photoproduction data is unjustified. The desire to make such interpretations should be restrained whilst development of reaction models continues and higher precision data are accumulated. One may anticipate that subsequent mutual feedback could lead to a well constrained phenomenology which may be employed in understanding $\gamma+p\to V+p$ processes and forging a tight link between the associated data and proton properties.
Acknowledgements.
We are grateful to Z.-F. Cui (崔著钫), T.-S. H. Lee, V. I. Mokeev, I. Strakovsky and J.-J. Wu for discussions and suggestions. Work supported by: National Natural Science Foundation of China, grant no. 12135007.
Data Availability Statement Data will be made available on reasonable request. [Authors’ comment: All information necessary to reproduce the results described herein is contained in the material presented above.]
Code Availability Statement Code/software will be made available on reasonable request. [Authors’ comment: No additional remarks.]
Appendix A Tensor structure of the $\gamma\to V$ transition
The essentially dynamical component of the reaction model, viz. the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element, has the tensor structure apparent in Eq. (20). The tensors take the following form:
$\tau^{1}_{\mu\nu\alpha}(q,P)$ $=q_{\alpha}R_{\mu\nu}\,$, (28a)
$\tau^{2}_{\mu\nu\alpha}(q,P)$ $=P_{\alpha}R_{\mu\nu}\,$, (28b)
$\tau^{3}_{\mu\nu\alpha}(q,P)$ $=q_{\alpha}q_{\nu}S_{\nu}\,$, (28c)
$\tau^{4}_{\mu\nu\alpha}(q,P)$ $=P_{\alpha}q_{\mu}S_{\nu}\,$, (28d)
$\tau^{5}_{\mu\nu\alpha}(q,P)$ $=q_{\alpha}P_{\mu}S_{\nu}\,$, (28e)
$\tau^{6}_{\mu\nu\alpha}(q,P)$ $=P_{\mu}T_{\alpha\nu}\,$, (28f)
$\tau^{7}_{\mu\nu\alpha}(q,P)$ $=q_{\mu}T_{\alpha\nu}\,$, (28g)
$\tau^{8}_{\mu\nu\alpha}(q,P)$ $=\delta_{\mu\alpha}S_{\nu}\,$, (28h)
$\tau^{9}_{\mu\nu\alpha}(q,P)$ $=P_{\alpha}P_{\mu}S_{\nu}\,$, (28i)
where
$R_{\mu\nu}$ $=\delta_{\mu\nu}-P_{\mu}P_{\nu}/P^{2}\,$, (29a)
$S_{\nu}$ $=q_{\nu}-P_{\nu}P\cdot q/P^{2}\,.$ (29b)

### v2

@@
-Preprint no. NJU-INP 105-25 Exclusive photoproduction of light and heavy vector mesons: thresholds to very high energies
+Preprint no. NJU-INP 105/25 Exclusive photoproduction of light and heavy vector mesons: thresholds to very high energies
 Abstract
@@
 Appealing to notions of vector meson dominance, the latter have been interpreted as vector-meson–proton scattering lengths.
-The body of results indicate that it is premature to link any $\gamma+p\to V+p$ data with, for instance, in-proton gluon distributions, the quantum chromodynamics trace anomaly, or pentaquark production.
+The body of results indicates that it is premature to link any $\gamma+p\to V+p$ data with, for instance, in-proton gluon distributions, the quantum chromodynamics trace anomaly, or pentaquark production.
 Further developments in reaction theory and higher precision data are required before the validity of any such links can be assessed.
@@
 provides a clean environment to study diffractive scattering from the proton and thereby obtain information about the reaction mechanism and both hadrons involved in the process.
+(Herein, we consider only ground-state vector-mesons.)
 Understanding the energy dependence of these cross sections has long challenged theory.
@@
 It is typically supposed to represent the collective exchange of a family of colourless, crossing-even states, with two-gluon exchange being the leading contributor Forshaw and Ross (2011).
-Connections with glueball physics are also discussed Meyer and Teper (2005); Boschi-Filho et al.
+Some phenomenological approaches distinguish between a hard (not necessarily perturbative) and soft Pomeron Donnachie and Landshoff (1998); and connections with glueball physics are also discussed Meyer and Teper (2005); Boschi-Filho et al.
 (2006).
@@
 3.
-Kinematics: $s=W^{2}=-(p_{2}^{2}+P^{2})$; $t=-(p_{2}-p_{1})^{2}$.
-With the high-energy behaviour of $\gamma+p\rightarrow V+p$ described by $\mathbb{P}$ exchange, it is natural to seek an explanation for the cross section at low-energies and determine whether a unified approach to the entire $s$-domain is possible.
+Kinematics: $s=W^{2}=-(p_{2}+P)^{2}$; $t=-(p_{2}-p_{1})^{2}$.
+Twenty and more years ago, a variety of such models were proposed as motivations for the measurement of $\gamma+p\rightarrow V+p$ photoproduction reactions; see, e.g., Refs.
+Nemchik et al.
+(1994); Cudell and Royen (1997); Frankfurt et al.
+(1999); Brodsky et al.
+(2001); Forshaw et al.
+(2004).
+A number were built upon the assumption of factorisation for hard exclusive processes, e.g., Nemchik et al.
+(1994); Frankfurt et al.
+(1999); Brodsky et al.
+(2001), and different studies canvassed the possibilities of distinct microscopic interaction mechanisms.
+Notwithstanding that, all such models attempted to draw a connection between $V$-meson photoproduction reactions and features of the target and produced meson, like in-target gluon structure functions, with such links drawn using Ansätze for the light-front wave functions of the bound states involved.
+Infrared freezing of the QCD running coupling was also sometimes assumed in the interaction models Nemchik et al.
+(1994); Cudell and Royen (1997), although with saturation values chosen which are significantly less than that associated with QCD’s process-independent effective charge Cui et al.
+(2020); Deur et al.
+(2024); Brodsky et al.
+(2024).
+Owing to their formulation assumptions, the models were only considered plausible at energies far above threshold.
+As will become clear, our framework represents a significant departure from such approaches.
+For instance, inter alia: no call is made to factorisation for hard exclusive processes; the required bound-state wave functions and constituent propagators are consistently calculated using an interaction that derives from modern analyses of QCD’s gauge sector; and predictions are simultaneously made for all $V$-meson final states, at energies that range, in each case, from threshold to very far above.
+Continuing with this last point, having noted that the high-energy behaviour of $\gamma+p\rightarrow V+p$ is described by $\mathbb{P}$ exchange, it is natural, as herein, to seek an explanation for the cross section at low-energies and determine whether a unified approach to the entire $s$-domain is possible.
 Both these issues were addressed in Ref.
@@
 2.C); and $\Gamma_{\mu}^{\gamma}$ is the associated dressed photon-quark vertex.
-Since we are focusing on photoproduction, then, with complete generality, one may use the Ball-Chiu form for $\Gamma_{\mu}^{\gamma}$ Ball and Chiu (1980).
+Since we are focusing on photoproduction, then, with complete generality, one may use the Ball-Chiu form for $\Gamma_{\mu}^{\gamma}$ Ball and Chiu (1980):
+$i\Gamma_{\mu}^{\rm BC}$ $(k_{1},k_{2})=\Sigma_{A}(k_{1}^{2},k_{2}^{2}))i\gamma_{\mu}+(k_{1}+k_{2})_{\mu}$
+$\times[\Delta_{A}(k_{1}^{2},k_{2}^{2})i\gamma\cdot(k_{1}+k_{2})+\Delta_{B}(k_{1}^{2},k_{2}^{2})]\,$, (5)
+where $\Sigma_{F}(k_{1}^{2},k_{2}^{2})=[F(k_{1})^{2}+F(k_{2}^{2})]/2$, $\Delta_{F}(k_{1}^{2},k_{2}^{2}))=[F(k_{1})^{2}-F(k_{2}^{2})]/[k_{1}^{2}-k_{2}^{2}]$, $F=A,B$.
+This Ansatz is complete for photoproduction because, with $Q^{2}=0$, it delivers the Ward identity result: $i\Gamma_{\mu}^{\rm BC}(k_{-},k_{-})=\partial_{\mu}S(k_{-})$.
 The remaining element in Eq.
-(4) is $\Gamma^{\mathbb{P}}_{\alpha}=\gamma_{\alpha}\beta_{\mathpzc q}$, the Pomeron-quark vertex.
-Here, $\beta_{\mathpzc q}$ is a constant strength factor and $\gamma_{\alpha}$ is a standard Dirac matrix.
-We have verified that replacing $\gamma_{\alpha}$ by some more sophisticated structure, such as the Ball-Chiu vertex Ball and Chiu (1980), for instance, has no material influence: any impact can be absorbed into a modest modification of $\beta_{\mathpzc q}$.
+(4) is $\Gamma^{\mathbb{P}}_{\alpha}$, the Pomeron-quark vertex.
+Following Refs.
+Donnachie and Landshoff (1984); Pichowsky and Lee (1997) and for the reasons detailed therein, we employ $\Gamma^{\mathbb{P}}_{\alpha}=\gamma_{\alpha}\beta_{\mathpzc q}$, where $\beta_{\mathpzc q}$ is a constant strength factor and $\gamma_{\alpha}$ is a standard Dirac matrix.
+[…]
 Working in the center of three-momentum frame of the vector-meson–proton system, the relevant four-momenta are:
@@
 The differential cross section for exclusive vector meson photoproduction can now be expressed:
-$\frac{d\sigma}{d\Omega}=\frac{1}{4\pi^{2}}\frac{m_{p}}{4W}\frac{|\vec{P}|}{k_{W}}\sum_{\text{proton spin}}(|{\cal I}_{1}|^{2}+|{\cal I}_{2}|^{2})\,$, (7)
+[…]
 where
@@
 Collins (1977); Irving and Worden (1977).
-These parameters depend on whether the final-state vector meson is light or heavy Lee et al.
-(2022).
-Table 1: Quark current-masses and parameters characterising the CSM interaction kernels for the mesons considered herein.
+As observed, e.g., in Refs.
+Donnachie and Landshoff (1998); Pichowsky and Lee (1997); Lee et al.
+(2022), data suggests that in comparing photoproduction reactions involving light or heavy vector-meson final states, different values of the pair $(\alpha_{0},\alpha_{1})$ are required; see, also, below.
+This observation encouraged some practitioners to suggest the existence of two Pomerons Donnachie and Landshoff (1998), one hard and one soft, with both emerging as the result of some (unknown) nonperturbative dynamics.
+Given the current QCD understanding of Pomeron phenomenology and structure, we elect not to express an opinion on this matter.
+Instead, in the reaction model advocated herein, there is a single (soft) exchanged object.
+The object fits within Pomeron-based schemes because, amongst other things, it provides a nonlocal color-singlet exchange that produces cross sections which rise with energy.
+This Pomeron represents essentially nonperturbative aspects of reactions involving hadrons; and the proton + Pomeron + quark interaction possesses an energy dependence that differs between light-quark meson and heavy-quark meson final states.
+It is, perhaps, worth noting that scale-dependent Pomeron trajectory parameters may be natural in some approaches, e.g., models based on AdS/CFT duality Dosch and Ferreira (2015).
+Table 1: All parameters used herein.
+Quark current-masses and parameters characterising the CSM interaction kernels for the mesons considered herein.
 The quark current masses are quoted at a renormalisation scale $\zeta=\zeta_{19}=19\,$GeV, i.e., in the far ultraviolet so as to avoid truncation ambiguity.
@@
 (2024).
-The same Pomeron trajectory is used for both $\rho^{0}$ and $\phi$ photoproduction Pichowsky and Lee (1997), and another for $J/\psi$ and $\Upsilon$ (Chekanov et al., 2002, ZEUS).
+The same Pomeron trajectory is used for both $\rho^{0}$ and $\phi$ photoproduction (Pichowsky and Lee, 1997, Sec.
+III.A), and another for $J/\psi$ and $\Upsilon$ (Chekanov et al., 2002, ZEUS).
 Pomeron-quark couplings, $\beta_{\mathpzc q}$, for both reaction models considered herein – Sects.
@@
 $\Upsilon_{b}$ 3.59 $0.6^{3}$ 0.8 5 0.36 1.2 0.115 0.016 0.638
+Figure 2: $\mathbb{P}$-am reaction model for $\gamma+p\to V+p$.
+Compared with our dynamical model, $\mathbb{P}$-dyn, sketched in Fig.
+1, $\mathbb{P}$-am replaces the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element by a momentum-independent $\gamma\to V$ transition coupling (akin to vector meson dominance) along with a form factor, $F_{V}$, that expresses a Pomeron–vector-meson form factor.
 2.3 VMD-like treatment of photon to meson transition
@@
 Sakinah et al.
-(2024) model; setting $v_{cN}=0$, so the quark-loop integration, based on a constituent quark model wavefunction, and VNV-N final state interactions (FSIs) do not contribute to the amplitude; and updating the Pomeron trajectories and Pomeron-quark couplings.
-Figure 2: $\mathbb{P}$-am reaction model for $\gamma+p\to V+p$.
-Compared with our dynamical model, $\mathbb{P}$-dyn, sketched in Fig.
-1, $\mathbb{P}$-am replaces the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$ transition matrix element by a momentum-independent $\gamma\to V$ transition coupling (akin to vector meson dominance) along with a form factor, $F_{V}$, that expresses a Pomeron–vector-meson form factor.
-The associated matrix element for diffractive vector meson photoproduction is
+(2024) model; setting $v_{cN}=0$, so the quark-loop integration, based on a constituent quark model wavefunction, and VNV-N final-state interactions (FSIs) do not contribute to the amplitude; and updating the Pomeron trajectories and Pomeron-quark couplings.
+The associated matrix element for diffractive vector meson photoproduction is (Sakinah et al., 2024, Eqs. (41-43)):
 ${\cal I}_{\mu}(W,t)\propto$ $\,\bar{u}(p_{2})\gamma_{\alpha}u(p_{1})\epsilon^{\lambda}_{\nu}(P)\Gamma_{\alpha,\mu\nu}G_{\mathbb{P}}(w^{2},t)$
 $\times[\beta_{\mathpzc q}F_{V}(t)][\beta_{l}F_{0}(t)]\,$, (14)
-with the tensor structure of the $\mathbb{P}$-$V$ vertex being Titov et al. (1998)
+with the tensor structure of the $\mathbb{P}$-$V$ vertex being a minimal Ansatz that ensures physical constraints (Titov et al., 1998, Sec. III.A), viz.
 $\Gamma_{\alpha,\mu\nu}=q_{\alpha}\delta_{\mu\nu}-q_{\mu}\delta_{\alpha\nu}\,.$ (15)
@@
 (3), $r,s,t,u$ represent colour, spinor, flavour matrix indices (as necessary).
-A realistic form of $\tilde{\mathpzc G}(y)$ is explained in Refs. Qin et al. (2011); Binosi et al. (2015):
+A realistic form of $\tilde{\mathpzc G}(y)$ is explained in Refs.
+Qin et al.
+(2011); Binosi et al.
+(2015):
 $\tilde{\mathpzc G}(y)$ $=\frac{8\pi^{2}}{\omega^{4}}De^{-y/\omega^{2}}+\frac{8\pi^{2}\gamma_{m}\mathcal{F}(y)}{\ln\big[\tau+(1+y/\Lambda_{\rm QCD}^{2})^{2}\big]}\,$, (18)
 […]
+At this point, with the quark current masses and parameter values in Table 1 left-block, the necessary gap and Bethe-Salpeter equations can be solved using standard algorithms Maris and Roberts (1997); Maris and Tandy (2006); Krassnigg (2008) to obtain each element in Eq.
+(4) and therewith $t_{\mu\alpha\nu}^{V}(q,P)$ itself.
+In solving all relevant Schwinger function equations, we use a mass-independent (chiral-limit) momentum-subtraction renormalisation scheme Chang et al.
+(2009), with renormalisation scale $\zeta=19\,$GeV.
+At this scale, the quark wave function renormalisation constants are practically unity.
 Applications to numerous systems and reactions have established Ding et al.
@@
 […]
-$\frac{d\sigma^{\gamma p\to Vp}}{d\Omega}\bigg|_{\text{th}}$ $=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{1}{64\pi}|T^{\gamma p\to Vp}|^{2},$ (18)
-$=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{\pi\alpha}{g_{V}^{2}}\frac{d\sigma^{Vp\to Vp}}{d\Omega}\bigg|_{\text{th}}$
-$=$ $\frac{|\vec{P}|}{|\vec{q}|}\frac{\pi\alpha}{g_{V}^{2}}|\alpha_{Vp}|^{2},$
-where $\alpha$ is the fine-structure constant, $\vec{q}$ and $\vec{P}$ are the initial and final momenta of the photon and vector meson, respectively – Eq.(5), and $\alpha_{Vp}$ is the model meson-proton scattering length.
+$\frac{d\sigma^{\gamma p\to Vp}}{d\Omega}\bigg|_{\text{th}}$ $=$ $\frac{|\vec{P}|}{k_{W}}\frac{1}{64\pi}|T^{\gamma p\to Vp}|^{2},$ (19)
+$=$ $\frac{|\vec{P}|}{k_{W}}\frac{\pi\alpha}{g_{V}^{2}}\frac{d\sigma^{Vp\to Vp}}{d\Omega}\bigg|_{\text{th}}$
+$=$ $\frac{|\vec{P}|}{k_{W}}\frac{\pi\alpha}{g_{V}^{2}}|\alpha_{Vp}|^{2},$
+where $\alpha$ is the fine-structure constant, $\vec{P}$ is the final momentum of the vector meson – Eq.(6), and $\alpha_{Vp}$ is the model meson-proton scattering length.
 The VMD ingredient is the coupling $g_{V}$, which is defined by the vector meson’s electromagnetic decay width $\Gamma_{V\to e^{+}e^{-}}$:
 $g_{V}^{2}=\frac{\pi\alpha^{2}m_{V}}{3\Gamma_{V\to\text{e}^{+}\text{e}^{-}}}.$ (20)
-Our prediction for this differential cross section is given in Eq.
+Suppose now that $V$-meson production near threshold is purely an $S$-wave process, then $\sigma_{\rm th}=4\pi d\sigma/d\Omega|_{\rm th}$.
+Then if one has precise total cross section data in hand, very close to threshold, it is possible to fit that data using a polynomial in $|\vec{P}|$, viz.
+write Strakovsky et al.
+(2020a)
+$\sigma_{\rm th}={\mathpzc a}_{1}|\vec{P}|+{\mathpzc a}_{3}|\vec{P}|^{3}+{\mathpzc a}_{5}|\vec{P}|^{5},$ (21)
+whereafter one has
+$|\alpha_{Vp}|^{2}={\mathpzc a}_{1}\frac{g_{V}^{2}k_{W}}{4\pi^{2}\alpha}\,.$ (22)
+Our prediction for the differential cross section is given in Eq.
 (8).
 That is the formula we use to evaluate all results.
-However, by equating our predictions with Eq.
-(18) one can arrive at model-specific VMD values for the meson-proton scattering length.
+However, by equating Eq.
+(8) with Eq.
+(19) one can arrive at model-specific VMD values for the meson-proton scattering length:
+$|\alpha_{Vp}|^{2}=\frac{1}{16\pi^{2}}\frac{m_{p}}{W}\frac{g_{V}^{2}}{\pi\alpha}\,\tfrac{1}{2}\!\!\!\sum_{\text{proton spin}}\!\!\!\left.(|{\cal I}_{1}|^{2}+|{\cal I}_{2}|^{2})\right|_{\rm th}\,.$ (23)
 These can be used as single parameters with which to evaluate the veracity of existing such model dependent inferences of meson-proton scattering lengths Strakovsky et al.
@@
 Then, the value has objective utility as a comparison between theory and experiment.
+We stress that one must use data very close to threshold if the fitted value of $\mathpzc a_{1}$ in Eq.
+(21) is to deliver a reliable empirical result for the slope parameter in Eq.
+(23).
 5 Results: cross sections
@@
 (3), (4).
-Accounting for current-conservation and the fact that $P_{\nu}\Gamma_{\nu}^{V}=0$, then, in all cases, $t_{\mu\alpha\nu}^{V}$ involves $9$ independent scalar functions which modulate the strength of the allowed tensor structures:
+Accounting for current conservation and the fact that $P_{\nu}\Gamma_{\nu}^{V}=0$, then, in all cases, $t_{\mu\alpha\nu}^{V}$ involves $9$ independent scalar functions which modulate the strength of the allowed tensor structures:
 $t_{\mu\nu\alpha}^{V}(q,P)=\sum_{i=1}^{9}F_{i}^{V}(q,P)\tau^{i}_{\mu\nu\alpha}(q,P).$ (24)
 The nine tensors are listed explicitly in A.
+Figure 4: $W$-dependence of total cross section for $\rho^{0}$ photoproduction, $W_{\rm th}^{\rho}=1.714\,$GeV.
+Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am.
+Data sources: black circles – (Struczinski et al., 1976, AHHM); cyan diamonds (fixed target) – (Aston et al., 1982, BCEGLMOPRS); green squares (Park et al., 1972; Ballam et al., 1972, SLAC); blue stars (Egloff et al., 1979, FNAL); red left and right triangles (Aid et al., 1996a; Andreev et al., 2020, H1); and red open squares – pseudodata inferred Strakovsky et al.
+(2025) from differential $\pi^{+}\pi^{-}p$ photoproduction cross sections measured with CLAS at JLab and analysed within the JM model Mokeev et al.
+(2016) (see text).
 Each scalar function in Eq. (24) is readily extracted using sensibly chosen projection operators. In practice, we find that the scalar function associated with the tensor structure
@@
 It is worth noting here that in fixed-target experiments, the photon energy is often used as the primary variable, viz. $E_{\gamma}(W)=k_{W}$ in Eq. (9).
-5.1 $\rho$ photoproduction
-Calculated using the Pomeron trajectory and quark–Pomeron couplings given in Table 1 – Row 1, our predictions for $\gamma+p\to\rho^{0}+p$ photoproduction are presented in Figs. 4 and 5.
-Figure 4: $W$-dependence of total cross section for $\rho^{0}$ photoproduction, $W_{\rm th}^{\rho}=1.714\,$GeV.
-Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am.
-Data sources: black circles – (Struczinski et al., 1976, AHHM); cyan diamonds (fixed target) – (Aston et al., 1982, BCEGLMOPRS); green squares (Park et al., 1972; Ballam et al., 1972, SLAC); blue stars (Egloff et al., 1979, FNAL); red left and right triangles (Aid et al., 1996a; Andreev et al., 2020, H1); and red open squares – pseudodata inferred Strakovsky et al.
-(2025) from differential $\pi^{+}\pi^{-}p$ photoproduction cross sections measured with CLAS at JLab and analysed within the JM model Mokeev et al.
-(2016) (see text).
 Figure 5: Differential cross section for $\rho^{0}$ photoproduction compared with data (Derrick et al., 1995a, ZEUS) (blue squares) in the range $60<W<$ $80$ GeV.
@@
 Pomeron trajectory and quark–Pomeron couplings used are given in Table 1 – Row 1.
+5.1 $\rho$ photoproduction
+Calculated using the Pomeron trajectory and quark–Pomeron couplings given in Table 1 – Row 1, our predictions for $\gamma+p\to\rho^{0}+p$ photoproduction are presented in Figs. 4 and 5.
 In Fig.
@@
 Such contributions decay rapidly with $W$, so that the diffractive $\mathbb{P}$-dyn reaction model becomes valid on $W\gtrsim 5\,$GeV.
+Figure 6: $W$-dependence of total cross section for $\phi$ photoproduction, $W_{\rm th}^{\phi}=1.958\,$GeV.
+Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am.
+Data: blue stars – (Egloff et al., 1979, FNAL); cyan triangles – (Busenitz et al., 1989, E401); black circles – (Derrick et al., 1996, ZEUS); and green open squares “pseudodata” Strakovsky et al.
+(2020b), obtained by interpolating CLAS differential cross sections (Dey et al., 2014, CLAS) as described in the text.
 Figure 5 depicts the differential cross section.
@@
 Regarding $\omega$-meson photoproduction, owing to its comparable structure, similar features and explanations pertain to the cross sections Yu et al. (2019).
-Figure 6: $W$-dependence of total cross section for $\phi$ photoproduction, $W_{\rm th}^{\phi}=1.958\,$GeV.
+Figure 7: $E_{\gamma}$ dependence of differential cross section for $\phi$ photoproduction, evaluated at $|t|=|t|_{\text{min}}$, Eq.
+(7), corresponding to forward-angle scattering.
 Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am.
-Data: blue stars – (Egloff et al., 1979, FNAL); cyan triangles – (Busenitz et al., 1989, E401); black circles – (Derrick et al., 1996, ZEUS); and green open squares “pseudodata” Strakovsky et al.
-(2020b), obtained by interpolating CLAS differential cross sections (Dey et al., 2014, CLAS) as described in the text.
+Data: black stars – (Ballam et al., 1973, SLAC); black up triangles – (Besch et al., 1974, Bonn); cyan squares – (Behrend et al., 1978, DESY); green – (Barber et al., 1982, Daresbury); black down triangles – (Barth et al., 2003, SAPHIR); blue open squares – (Mibe et al., 2005, LEPS); red circles – (Seraydaryan et al., 2014, CLAS).
 5.2 $\phi$ photoproduction
@@
 (2020b) generated pseudodata by interpolating differential cross sections measured at JLab (Dey et al., 2014, CLAS) and integrating over the differential solid angle.
-The $\mathbb{P}$-dyn prediction aligns well with this pseudodata, too; although it misses the “bump” apparent in the pseudodata.
+The $\mathbb{P}$-dyn prediction aligns well with this pseudodata, too; although it misses the “bump” that inspection reveals in the pseudodata (Strakovsky et al., 2020b, Fig.
+1).
 We note that this bump may be merely an artefact of the pseudodata construction procedure.
-Figure 7: $E_{\gamma}$ dependence of differential cross section for $\phi$ photoproduction, evaluated at $|t|=|t|_{\text{min}}$, Eq.
-(6), corresponding to forward-angle scattering.
-Solid purple curve – $\mathbb{P}$-dyn; dashed black curve – $\mathbb{P}$-am.
-Data: black stars – (Ballam et al., 1973, SLAC); black up triangles – (Besch et al., 1974, Bonn); cyan squares – (Behrend et al., 1978, DESY); green – (Barber et al., 1982, Daresbury); black down triangles – (Barth et al., 2003, SAPHIR); blue open squares – (Mibe et al., 2005, LEPS); red circles – (Seraydaryan et al., 2014, CLAS).
 The $E_{\gamma}$ dependence of the forward-angle $\phi$ photoproduction differential cross section is drawn in Fig.
@@
 Compared with data, the $\mathbb{P}$-dyn reaction model delivers a better global description than $\mathbb{P}$-am; hence, our dynamical treatment of the $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to\phi$ transition is once more seen to be important.
+It should nevertheless be remarked that some structure seems apparent in Fig.
+7 on a domain $E_{\gamma}\simeq 2.5\,$GeV.
+This is reproduced by neither our reaction model nor an earlier analysis Titov and Lee (2003) that also included $s$-channel resonance contributions with subsequent $\phi$-meson radiation.
+Since hadron interactions in these channels are poorly known, it might be that the structure owes, at least in part, to singularities associated, e.g., with meson + baryon FSIs Ryu et al.
+(2014).
+All that can now be safely concluded is that diffractive mechanisms do not reproduce $E_{\gamma}\simeq 2.5\,$GeV structure, and more data are required before it may properly be understood.
 A
@@
 These observations show that whilst modern Poincaré covariant calculations do indicate that the proton mass radius, $r_{m}$, is smaller than its charge radius Yao et al.
+(2025); Binosi et al.
 (2026), owing to the different character of the effective probe Xu et al.
@@
 (Adhikari et al., 2023, GlueX) data suggest a role for some nondiffractive contributions to the reaction on $|t|\gtrsim 3\,$GeV2.
-This domain is not reached by the Ref.
+This possibility may also receive support from the improved match between the Ref.
+Sakinah et al.
+(2024) (FSI) model and data on that domain.
+The domain $|t|\gtrsim 3\,$GeV2 is not reached by the Ref.
 (Duran et al., 2023, $J/\psi$-007) data: on their domain of coverage, those data are entirely consistent with our $\mathbb{P}$-dyn reaction model.
@@
 1 is crucial in achieving agreement with data: the $\mathbb{P}$-am model fails on this domain, becoming viable only on $W\gtrsim 5\,$GeV.
+A
+B
+Figure 16: $|t|$ dependence of differential cross section for $\Upsilon$ photoproduction.
+Panel A.
+W=11W=11\,GeV.
+Panel B.
+$\mathbb{P}$-dyn: solid purple – $W=100\,{\rm GeV}$; long-dashed purple – 500500\,GeV; dot-dashed purple – 11\,TeV.
+$\mathbb{P}$-am: dashed black – $W=100\,{\rm GeV}$; dot-dot-dashed black – 500500\,GeV; dotted black – 11\,TeV.
 Emulating Ref.
@@
 These facts further impair attempts to connect near threshold $J/\psi$ photoproduction data with the in-proton gluon GPD.
-A
-B
-Figure 16: $|t|$ dependence of differential cross section for $\Upsilon$ photoproduction.
-Panel A.
-W=11W=11\,GeV.
-Panel B.
-$\mathbb{P}$-dyn: solid purple – $W=100\,{\rm GeV}$; long-dashed purple – 500500\,GeV; dot-dashed purple – 11\,TeV.
-$\mathbb{P}$-am: dashed black – $W=100\,{\rm GeV}$; dot-dot-dashed black – 500500\,GeV; dotted black – 11\,TeV.
 On the other hand, the reaction model in Ref.
@@
 Table 4: VMD-based scattering lengths – Eq.
-(18) – inferred from near-threshold $V$-meson photoproduction differential cross sections calculated using the $\mathbb{P}$-dyn and $\mathbb{P}$-am reaction models.
+(23) – obtained from near-threshold $V$-meson photoproduction differential cross sections calculated using the $\mathbb{P}$-dyn and $\mathbb{P}$-am reaction models.
 $q_{\rm th}$ is the photon momentum at threshold in the centre-of-mass frame.
-The value of $|\alpha_{\Upsilon p}^{\rm VMD}|$ was estimated in Ref.
+The values of $|\alpha_{Vp}^{\rm VMD}|$ in the final column were calculated using Eq.
+(22) and values of ${\mathpzc a}_{1}$ inferred from Refs.
+Strakovsky et al.
+(2025, 2020b, 2020a, 2021).
+Note that for $|\alpha_{\Upsilon p}^{\rm VMD}|$ the value of ${\mathpzc a}_{1}$ in Eq.
+(21) was estimated in Ref.
 Strakovsky et al.
@@
 (2021), which is an earlier version of the GPD model discussed above Guo et al.
-(2023): the GPD reaction model delivers results that disagree with ours by an order of magnitude and more.
+(2023): the GPD reaction model delivers results that disagree with ours by an order of magnitude.
 $V$ $m_{V}$/GeV $q_{th}$/GeV $\Gamma_{V\to e^{+}e^{-}}$/keV $|\alpha_{Vp}^{\rm dyn}|$/fm $|\alpha_{Vp}^{\rm am}|$/fm $|\alpha_{Vp}^{\rm VMD}|$/fm
-$\rho^{0}\$0.775 0.6 6.97 0.136 0.072 $0.23\pm 0.03$ Strakovsky et al. (2025)
-$\phi^{0}\$1.019 0.75 1.27 0.087 0.048 $0.063\pm 0.010$ Strakovsky et al. (2020b)
-$J/\psi\$3.097 1.91 5.53 $3.08\times 10^{-3}$ $1.88\times 10^{-3}\$$3.08(55)\times 10^{-3}$ Strakovsky et al. (2020a)
-$\Upsilon\$9.46 5.16 1.34 $5.87\times 10^{-5}$ $1.86\times 10^{-5}\$$0.51(3)\times 10^{-3}$ Strakovsky et al. (2021)
+$\rho^{0}\$0.775 0.6 6.97 0.184 0.097 $0.31(04)$
+$\phi^{0}\$1.019 0.75 1.27 0.123 0.069 $0.090(14)$
+$J/\psi\$3.097 1.91 5.53 $6.34\times 10^{-3}$ $3.89\times 10^{-3}\$$6.4(1.1)\times 10^{-3}$
+$\Upsilon\$9.46 5.16 1.34 $196\times 10^{-3} 0.0$$062\times 10^{-3}\$$1.7(0.1)\times 10^{-3}$
 In the standard Pomeron-alone framework, the exponent should be Donnachie and Landshoff (1992):
@@
 (8) and this is the formula we used above to calculate all results.
-Figure 19: VMD-based scattering lengths – Eq. (18) – inferred from near-threshold vector meson photoproduction differential cross sections. Pictorial representation of the results listed in Table 4.
+Figure 19: VMD-based scattering lengths – Eq.
+(23) – calculated from near-threshold vector meson photoproduction differential cross sections.
+Pictorial representation of the results listed in Table 4.
+Legend.
+$\mathbb{P}$-dyn – black boxes; $\mathbb{P}$-am – red diamonds; empirical inferrences – blue stars.
 Here, for the sake of a comparison between our predictions and the model-specific VMD-values of meson-proton scattering lengths inferred in Refs.
@@
 (2020a, b, 2021, 2025), we equate our near-threshold results with Eq.
-(18) to arrive at a crude estimate of $|\alpha_{Vp}|$.
+(19) to arrive at an estimate of the slope parameter $|\alpha_{Vp}|$ via Eq.
+(23).
 The results are listed in Table 4 and plotted in Fig.
@@
 It is worth recording the following pattern in our predictions:
-$|\alpha_{\Upsilon p}|\ll|\alpha_{J/\psi p}|\ll|\alpha_{\phi p}|<|\alpha_{\rho^{0}p}|\,.$ (26)
+$|\alpha_{\Upsilon p}|<|\alpha_{J/\psi p}|\ll|\alpha_{\phi p}|<|\alpha_{\rho^{0}p}|\,.$ (30)
 The $\mathbb{P}$-dyn results may be described by
-$|\alpha_{Vp}|\approx{\mathpzc a}_{0}/m_{V}^{3},\quad{\mathpzc a}_{0}=(0.41\,{\rm GeV})^{3}\,{\rm fm}.$ (27)
+$|\alpha_{Vp}|\approx{\mathpzc a}_{0}/m_{V}^{2},\quad{\mathpzc a}_{0}=(0.34\,{\rm GeV})^{2}\,{\rm fm}.$ (31)
 As observed in Sect.
 4, one need not interpret $|\alpha_{Vp}|$ as the VpV-p scattering length; instead, it can simply be viewed as a near-threshold slope parameter, in which case the value has objective utility as a comparison between theory and experiment.
+We note here that we have compared results obtained using Eq.
+(23) with those obtained from our total cross sections using Eqs.
+(21), (22).
+The results are identical so long as one fits the total cross sections very near threshold, viz.
+on $|\vec{P}|\in[0,0.1]\,$GeV.
+If one is forced by (experimental) circumstances to begin further from threshold and use a larger range of momenta, then one cannot be certain that the estimate of ${\mathpzc a}_{1}$ is reliable.
 Regarding $V=\rho^{0}$, there is qualitative agreement between our near-threshold slope and that associated with available data.
@@
 Hence, the agreement with data is good, especially when one considers the scatter in available near threshold data; see Figs.
-7 – 10.
+7 – 10, and the use of such data in Ref.
+Strakovsky et al.
+(2020b) to infer a value of ${\mathpzc a}_{1}^{\phi}$.
 Owing to the heightened interest in $J/\psi$ photoproduction, a good deal of modern data is available near threshold.
@@
 In our view, given its success with $J/\psi$ photoproduction, the $\mathbb{P}$-dyn reaction model should be reliable for $\Upsilon$, too.
-However, its prediction is an order of magnitude and more smaller than that listed in Ref.
+However, its prediction is an order of magnitude smaller than that inferred from Ref.
 Strakovsky et al.
@@
 8 Summary and perspective
-We described and employed a reaction model, $\mathbb{P}$-dyn, for $\gamma+p\to V+p$ photoproduction, $V=\rho^{0},\phi,J/\psi,\Upsilon$, that exposes the quark-antiquark content of the dressed photon in making the transition $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$, where $\mathpzc q$} depends on $V$, and couples the intermediate ${\mathpzc q}\bar{\mathpzc q}$ system to the proton’s valence quarks via Pomeron exchange [Fig.
+We described and employed a reaction model, $\mathbb{P}$-dyn, for $\gamma+p\to V+p$ photoproduction, $V=\rho^{0},\phi,J/\psi,\Upsilon$, that exposes the quark+antiquark content of the dressed photon in making the transition $\gamma\to{\mathpzc q}\bar{\mathpzc q}+\mathbb{P}\to V$, where $\mathpzc q$} depends on $V$, and couples the intermediate ${\mathpzc q}\bar{\mathpzc q}$ system to the proton’s valence quarks via Pomeron exchange [Fig.
 1].
@@
 Acknowledgements.
-We are grateful to Z.-F.
-Cui (崔著钫), T.-S.
-H.
-Lee, V.
-I.
-Mokeev, I.
-Strakovsky and J.-J.
-Wu for discussions and suggestions.
-Work supported by: National Natural Science Foundation of China, grant no.
-12135007.
+We are grateful to T.-S. H. Lee, V. I. Mokeev, I. Strakovsky and J.-J. Wu for discussions and suggestions. Work supported by: National Natural Science Foundation of China, grant no. 12135007.
 Data Availability Statement Data will be made available on reasonable request.